### v2

Abstract.
We study $\eta$-Einstein Sasakian structures on Lie algebras, that is, Sasakian structures whose associated Ricci tensor satisfies an Einstein-like condition. We divide into the cases in which the Lie algebra’s centre is non-trivial (and necessarily one-dimensional) from those where it is zero. In the former case we show that any Sasakian structure on a unimodular Lie algebra is $\eta$-Einstein. As for centreless Sasakian Lie algebras, we devise a complete characterisation under certain dimensional assumptions regarding the action of the Reeb vector. Using this result, together with the theory of normal $j$-algebras and modifications of Hermitian Lie algebras, we construct new examples of $\eta$-Einstein Sasakian Lie algebras and solvmanifolds, and provide effective restrictions for their existence.
Key words and phrases:
Sasakian Lie algebra, $\eta$-Einstein metric, normal $j$-algebra, invariant structure, Kähler-exact Lie algebra
Dedicated to our dear Marisa Fernández
1. Introduction
A Sasakian manifold is the odd-dimensional counterpart of a Kähler manifold, due to the fact that a Riemannian manifold $(M,g)$ of dimension $2n+1$ is Sasakian if and only if its Riemannian cone $M\times\mathbb{R}^{+}$ is Kähler. Sasakian structures on $(M,g)$ are triples $(\varphi,\eta,\xi)$, consisting of an endomorphism $\varphi\in\operatorname{\mathfrak{so}}(TM)$, a vector field $\xi\in TM$ (the Reeb vector) and the metric-dual form $\eta\in TM^{*}$, satisfying certain compatibility conditions (§2.1). These manifolds have been studied at length since Sasaki introduced them $60$-odd years ago.
A distinguished family of Sasakian manifolds are those whose metric is Einstein, meaning that the associated Ricci tensor satisfies $\operatorname{Ric}=\lambda g$ for some $\lambda\in\mathbb{R}$. They are naturally called Sasaki-Einstein manifolds, and enjoy remarkable properties. For instance, it turns out that $\lambda=2n$, so $g$ has positive Ricci curvature, whence $\pi_{1}(M)$ is finite in case $g$ is complete, by Myers’ theorem. The prototypical examples are the spheres $S^{2n+1}$ with the standard Kähler metric. The textbooks [10] and [14] have become classical references on Sasakian geometry.
Okumura [32] defined a generalisation of Sasaki-Einstein manifolds by imposing a condition that relates $\operatorname{Ric}$ not only to $g$, but also to the Reeb $1$-form $\eta$. A Sasakian manifold is called $\eta$-Einstein if $\operatorname{Ric}=\lambda g+\nu\,\eta\otimes\eta$ for some constants $\lambda,\nu\in\mathbb{R}$. The Sasakian-Einstein case is recovered for $\nu=0$. The constants $\lambda$ and $\nu$ are not arbitrary as they must obey $\lambda+\nu=2n$, and it can be shown that demanding $g$ be $\eta$-Einstein is equivalent to saying the metric is Einstein on the distribution transversal to the Reeb vector. From that, every Sasakian $\eta$-Einstein manifold has constant scalar curvature equal to $2n(\lambda+1)$. It was shown in [15] that the value $\lambda=-2$ plays an important role in this context, for such manifolds can now be labelled positive (if $\lambda>-2$), null ($\lambda=-2$) and negative ($\lambda<-2$). In particular, therefore, Sasaki-Einstein manifolds are positive.
According to [25], a $3$-dimensional compact Sasakian manifold is diffeomorphic to the quotient of either $S^{3}\cong\operatorname{\operatorname{SU}}(2)$, the Heisenberg group $H_{3}$ or the universal cover $\widetilde{\operatorname{\operatorname{SL}}}(2,\mathbb{R})$ by some lattice (a co-compact discrete subgroup). These three model geometries correspond precisely to positive, null or negative Sasakian $\eta$-Einstein structures. That the universal cover of any of the above is a Lie group motivates the study of ($\eta$-Einstein) Sasakian structures on Lie groups. The investigation of Lie groups equipped with left-invariant Sasakian structures was initiated in [5]. That article explains why the Lie algebras of Sasakian Lie groups can be divided into two classes: the ones with one-dimensional centre (which has to be spanned by the Reeb vector), and the ones with trivial centre. The first class is well understood, for it consists of central extensions of Kähler Lie algebras by the cocycle given by the Kähler form. Much less is known regarding Sasakian Lie algebras with trivial centre. The same paper provides us with a classification of $5$-dimensional Sasakian Lie algebras, and highlights the $\eta$-Einstein ones among them: among the $5$-dimensional Lie algebras admitting an $\eta$-Einstein Sasakian structure are the Heisenberg algebra $\mathfrak{h}_{5}$, a very specific non-unimodular solvable Lie algebra and the product $\mathfrak{sl}(2,\mathbb{R})\times\mathfrak{aff}(\mathbb{R})$. As it turns out, these structures are all negative or null $\eta$-Einstein.
The present article’s goal is to study higher-dimensional Lie groups equipped with a left-invariant $\eta$-Einstein Sasakian structure. To do so the first thing is to split the two cases: non-trivial centre and trivial centre. Over the former class the analysis is simpler, and we are able to show (Proposition 3.3) that a Sasakian Lie algebra with non-trivial centre is $\eta$-Einstein (with Einstein constant $\lambda$) precisely if it is the central extension of a Kähler-Einstein Lie algebra (with defining constant $\lambda+2$). Moreover, Corollary 3.4 proves that if a Sasakian Lie algebra with non-trivial centre is unimodular then the structure is automatically $\eta$-Einstein, and actually null ($\lambda=-2$). In particular, Sasakian solvmanifolds whose Lie algebra has non-trivial centre must be null $\eta$-Einstein, and we provide in Example 3.5 original examples of solvmanifolds equipped with null $\eta$-Einstein Sasakian geometry.
At the opposite end, if the centre is trivial, the Reeb vector $\xi$ no longer commutes with everything, but we can use it to split the Lie algebra $\mathfrak{g}$ into an orthogonal sum of the kernel of $\operatorname{ad}_{\xi}$ (a Sasakian subalgebra with non-trivial centre) and the image of $\operatorname{ad}_{\xi}$. Alas, the computations in this setting are considerably more involved and give little insight per se. We therefore restrict to the case where the image of $\operatorname{ad}_{\xi}$ is two-dimensional, which at least geometrically is reasonable. This extra assumption enables us to characterise all centreless Sasakian Lie algebras of this type (Proposition 4.1 and Theorem 4.3). Our characterisation involves a Kähler Lie algebra $\mathfrak{h}_{1}$ (with exact Kähler form) and a distinguished vector $H_{0}$ in the commutator ideal of $\mathfrak{h}_{1}$. As it turns out, the element $H_{0}$ plays a key role in the study. For instance, if $H_{0}=0$ it follows that $\mathfrak{g}$ is never solvable, whereas if $H_{0}\neq 0$ and $\mathfrak{h}_{1}$ is solvable then $\mathfrak{g}$ must be solvable as well. Regarding the $\eta$-Einstein condition, we show in Proposition 4.10 that under the assumption $H_{0}=0$, the Sasakian Lie algebra $\mathfrak{g}$ is $\eta$-Einstein if and only if $\mathfrak{h}_{1}$ is Kähler-Einstein. The generic case ($H_{0}\neq 0$) is more difficult to address, and to tackle it we suppose $\mathfrak{h}_{1}$ is solvable. Since semisimple Lie algebras cannot carry Kähler forms (nor any symplectic form, by the way), this further hypothesis is rather natural in the light of the Levi decomposition. The same assumption becomes even more significant due to the non-accidental appearance of so-called normal $j$-algebras (or $J$-algebras according to [2]). These were introduced by Piatetskii-Shapiro and much studied by, first and foremost, Dorfmeister, D’Atri and Dotti. After a careful description, we use them in Theorem 5.16 to construct new examples of $\eta$-Einstein Sasakian Lie algebras of negative type.
Subsequently, we bring in the concept of modification of a metric Lie algebra with a two-fold purpose: first, we show in Theorem 6.7 that if $G$ is an arbitrary Lie group of dimension $\geqslant 5$ that admits a left-invariant $\eta$-Einstein Sasakian structure and contains a lattice, then $G$ is solvable and the $\eta$-Einstein structure is null. Secondly, we prove that any Kähler-exact Lie algebra is the modification of a normal $j$-algebra. We then employ this fact for showing that there are indeed only finitely many possibilities to pick $H_{0}$ (Proposition 7.5), based on which we also provide obstructions to the existence of $\eta$-Einstein structures (Proposition 7.7).
Finally, in Proposition 8.1 we consider the extreme case when the image of $\operatorname{ad}_{\xi}$ is largest possible. We prove that this can only happen in dimension $3$, so the algebras of the case are $\mathfrak{su}(2)$ or $\mathfrak{sl}(2,\mathbb{R})$.
Acknowledgements.
This work is part of Nuñez’ PhD thesis, and preliminary results were presented at the 2024 conference ‘Special holonomy and geometric structures on complex manifolds’ held at impa. He thanks the PhD programme at uff, and capes, for the financial support during his graduate studies and for the PDSE grant 6/2024.
AN and SGC are grateful to famaf (Facultad de Matemática, Astronomía, Física y Computación at Universidad Nacional de Córdoba) and ciem (Centro de Investigación y Estudios de Matemática of conicet) for the unwavering warm hospitality in Córdoba. The three authors are indebted to Anna Fino, Gueo Grantcharov, Giulia Dileo, Ilka Agricola and Jorge Lauret for insightful conversations and for pointing out useful references. They thank the referee for reading the paper’s first version carefully and making spot-on comments and improvements.
SGC is a member of indam (Istituto Nazionale di Alta Matematica ‘Francesco Severi’).
2. Preliminaries
2.1. Sasakian manifolds
We are interested in Sasakian Lie algebras and their curvature features, with an eye to $\eta$-Einstein metrics. For completeness we shall begin by laying out the ambient framework, the standard notation and swiftly review a few known results.
For starters, an almost contact structure on an odd-dimensional smooth real manifold $M$ is a triple $(\varphi,\eta,\xi)$ where $\xi$ is a nowhere-vanishing vector field (called the Reeb vector field), $\eta$ is a $1$-form satisfying $\eta(\xi)=1$ and $\varphi$ is an endomorphism of $TM$ such that
$\varphi^{2}=-\operatorname{\operatorname{Id}}+\xi\otimes\eta.$
One refers to $(M,\varphi,\eta,\xi)$ as an almost contact manifold. The above conditions force $\varphi(\xi)=0$ and $\eta\circ\varphi=0$, and as a consequence the tangent bundle pointwise splits into a horizontal hyperplane plus a line
$TM=\operatorname{Ker}\eta\oplus\mathbb{R}\xi,$
and furthermore $\operatorname{Ker}\eta=\operatorname{Im}\varphi$. Associated with $\varphi$ is the Nijenhuis $(2,1)$-tensor $N_{\varphi}$:
(1) $N_{\varphi}(X,Y):=\varphi^{2}[X,Y]+[\varphi X,\varphi Y]-\varphi[\varphi X,Y]-\varphi[X,\varphi Y],$
and the almost contact structure $(\varphi,\eta,\xi)$ is called normal if
$N_{\varphi}=-\textup{d}\eta\otimes\xi.$
The latter is equivalent [36] to requiring that the almost complex structure
(2) $J\left(X,f\tfrac{d}{dt}\right)=\left(\varphi X-f\xi,\eta(X)\tfrac{d}{dt}\right)$
on $M\times\mathbb{R}$ be integrable, where $f\in\mathcal{C}^{\infty}(M\times\mathbb{R})$ and $t$ is a coordinate on $\mathbb{R}$.
One says a Riemannian metric $g$ on an almost contact manifold $(M,\varphi,\eta,\xi)$ is compatible with the almost contact structure if
$g(\varphi X,\varphi Y)=g(X,Y)-\eta(X)\eta(Y)$
for any $X,Y\in TM$. This defines a reduction of the manifold’s structure group to $\operatorname{\operatorname{U}}(n)$ and $(\varphi,\eta,\xi,g)$ becomes known as an almost contact metric structure. In such a quadruple the tensor $\varphi$ is skew-symmetric for $g$, and the Reeb field is the metric dual of $\eta$. Having a metric and $\varphi$ enables one to define the so-called fundamental $2$-form $\Phi\in\bigwedge^{2}T^{*}M$ by
$\Phi(X,Y):=g(X,\varphi Y),$
in perfect analogy to the almost Hermitian setting. The classical Chinea–González classification of types of almost contact metric structure based on the intrinsic torsion has finally found a very geometric interpretation and reworking in [1].
As $M$ will typically be $(2n+1)$-dimensional, we will write $M^{2n+1}$ to remind ourselves.
Definition 2.1.
An almost contact metric structure $(\varphi,\eta,\xi,g)$ on a manifold $M^{2n+1}$ is called Sasakian provided it is normal and $\textup{d}\eta=2\Phi$ (making $\Phi$ exact). The data $(M,\varphi,\eta,\xi,g)$ is called a Sasakian manifold.
The relation $\textup{d}\eta=2\Phi$ implies $\eta$ is a contact form (ie, $\eta\wedge(d\eta)^{n}$ is a volume form).
There are a number of reasons for which Sasakian manifolds represent the most significant class of almost contact metric manifolds. First and foremost is their strong link to Kähler manifolds: the Riemannian cone of an almost contact metric manifold $M$, equipped with the structure induced by (2), is Kähler if and only if $M$ is Sasakian (see [13], but also [28]).
As Sasakian manifolds are central to the present work, we shall recall a few essential properties that will be useful in upcoming sections. They are standard and proofs can be found in [10, 14].
Henceforth $\nabla$ will indicate the Levi-Civita connection associated with $g$, and $R$ and $\operatorname{Ric}$ will be the Riemannian and Ricci tensors. Our convention for the Riemann curvature is $R_{x,y}:=[\nabla_{x},\nabla_{y}]-\nabla_{[x,y]}$.
Proposition 2.2.
For any vector fields $X,Y$ on a Sasakian manifold $(M^{2n+1},\varphi,\eta,\xi,g)$ we have:
  1. (1)
    $\nabla_{X}\xi=-\varphi X$  and  $\nabla_{\xi}X=[\xi,X]-\varphi X$,
  2. (2)
    $(\nabla_{X}\varphi)Y=g(X,Y)\xi-\eta(Y)X$,
  3. (3)
    $R_{X,Y}\xi=\eta(Y)X-\eta(X)Y$,
  4. (4)
    $\operatorname{Ric}(\xi,X)=2n\,\eta(X)$.
In particular, the Reeb vector is a Killing vector field ($\nabla_{\xi}\xi=0$) and $\operatorname{Ric}(\xi,\xi)=2n$, so a Sasakian metric is never Ricci-flat.
Our true focus is a distinguished family of Sasakian manifolds, called $\eta$-Einstein because they are defined in terms of the Ricci tensor just like ordinary Einstein manifolds are. They were introduced by Okumura [32] (see also [35]), and a good starting source of study is [15].
Definition 2.3.
A Sasakian manifold $(M^{2n+1},\varphi,\xi,\eta,g)$ is said to be $\bm{\eta}$-Einstein if the Ricci tensor satisfies the constraint
$\operatorname{Ric}=\lambda g+(2n-\lambda)\eta\otimes\eta$
for some constant $\lambda\in\mathbb{R}$. More specifically, it is called positive ($\eta$-Einstein) if $\lambda>-2$, null if $\lambda=-2$, and negative if $\lambda<-2$.
Observe that the $\eta$-Einstein condition reads
$\operatorname{Ric}(U,V)=\lambda g(U,V)$
for all Reeb-transversal vector fields $U,V$. Another immediate consequence of the definition is that every $\eta$-Einstein Sasakian manifold automatically has constant scalar curvature $2n(\lambda+1)$.
2.2. Sasakian Lie algebras
We shall study left-invariant structures on Lie groups $G$. As these are uniquely determined by their value at the group’s identity, we are entitled to consider structures on Lie algebras $\mathfrak{g}=T_{e}G$. Just for the purpose of collecting all the pieces of data presented earlier, let us recall that a metric Lie algebra $(\mathfrak{g},\langle\cdotp,\cdotp\rangle)$ is a Lie algebra $\mathfrak{g}$ equipped with a positive-definite, symmetric bilinear form $\langle\cdotp,\cdotp\rangle$. Any such inner product corresponds uniquely to a left-invariant Riemannian metric on each connected group $G$ integrating $\mathfrak{g}$.
The linear map $\textup{d}:\mathfrak{g}^{*}\longrightarrow\bigwedge^{2}\mathfrak{g}^{*}$ that sends a $1$-form $\alpha$ to $\textup{d}\alpha$,
$\textup{d}\alpha(x,y):=-\alpha[x,y],$
is the (Lie algebra) exterior derivative or Chevalley-Eilenberg differential. It extends to a graded linear map $\textup{d}:\bigwedge^{k}\mathfrak{g}^{*}\longrightarrow\bigwedge^{k+1}\mathfrak{g}^{*}$ in any degree which, spelt out, reads
$(\textup{d}\phi)(x_{0},x_{1},\ldots,x_{k})=\sum_{0\leqslant i<j\leqslant k}(-1)^{i+j}\phi([x_{i},x_{j}],x_{0},\ldots,\widehat{x_{i}},\dots,\widehat{x_{j}},\dots,x_{k})$
where the hat means that the term is omitted.
If $(e_{1},\ldots,e_{n})$ is a basis of $\mathfrak{g}$, we will denote by $(e^{1},\ldots,e^{n})$ the dual basis of $\mathfrak{g}^{*}$ and conventionally write $e^{ij}:=e^{i}\wedge e^{j}$. Then the string
$(\textup{d}e^{1},\textup{d}e^{2},\ldots,\textup{d}e^{n})$
encodes the structure equations $e_{i},e_{j}]=\sum_{k}c_{ij}^{k}e_{k}$ since $\textup{d}e^{k}=-\sum_{i<j}c_{ij}^{k}e^{ij}$. This is very practical, for the Jacobi identity on $\mathfrak{g}$ is equivalent to asking the exterior algebra $(\bigwedge\mathfrak{g}^{*},\textup{d})$ be a complex.
Two major players in our work are the $2$-dimensional affine Lie algebra and the $(2n+1)$-dimensional Heisenberg algebra, so we recall them for the record:
  • $\mathfrak{aff}(\mathbb{R})$ is spanned by vectors $e_{1},e_{2}$ with structure equations $(0,e^{21})$;
  • $\mathfrak{h}_{2n+1}=\operatorname{span}(e_{1},\ldots,e_{n},f_{1},\ldots,f_{n},z)$ with equations $(0,\ldots,0,f^{1}\wedge e^{1}+\cdots+f^{n}\wedge e^{n})$.
Definition 2.4.
A Sasakian Lie algebra consists of a metric Lie algebra $\big(\mathfrak{g},\langle\cdotp,\cdotp\rangle\big)$, a vector $\xi\in\mathfrak{g}$, its metric dual $\eta\in\mathfrak{g}^{*}$ and an endomorphism $\varphi\in\mathfrak{g}^{*}\otimes\mathfrak{g}$ subject to:
$\begin{array}[]{cc}\varphi^{2}=-\operatorname{\operatorname{Id}}+\xi\otimes\eta,\qquad\langle\varphi x,\varphi y\rangle=\langle x,y\rangle-\eta(x)\eta(y)\\[5.69054pt] \textup{d}\eta(x,y)=2\langle x,\varphi y\rangle,\qquad N_{\varphi}=-d\eta\otimes\xi,\end{array}$
with $N_{\varphi}$ as per (1).
As before, we set $\Phi(x,y):=\langle x,\varphi y\rangle$. Now, the normality condition is equivalent (see [4, Proposition 3.1]) to the pair of requirements
$\operatorname{ad}_{\xi}\circ\varphi=\varphi\circ\operatorname{ad}_{\xi},\qquad[\varphi x,\varphi y]-[x,y]=\varphi\big([\varphi x,y]+[x,\varphi y]\big)$
for all $x,y\in\operatorname{Ker}\eta$.
Notation-wise, let us name the nullspace of $\eta$ and the projection of the Lie bracket to this nullspace by
[…]
where $\operatorname{pr}_{\mathfrak{h}}:\mathfrak{g}\to\mathfrak{h}$ denotes the projection coming from the decomposition $\mathfrak{g}=\mathbb{R}\xi\oplus\mathfrak{h}$.
The next two results, proved in [5], contain the starting point of our investigations. First, the centre $\mathfrak{z}$ of a Sasakian Lie algebra is either trivial or spanned by $\xi$ [cit, proposition 3.4]. We shall say that $\mathfrak{g}$ has a centre or has no centre/is centreless in order to distinguish the two situations. Secondly, if the Reeb vector is central then $\theta$ is a Lie bracket on $\mathfrak{h}$, and we mod $\xi$ out so that the algebra $(\mathfrak{h}\cong\mathfrak{g}/\mathbb{R},\theta$) acquires a Kähler structure
[…]
induced from the geometry of $\mathfrak{g}$. The same recipe, read backwards, prescribes how the central extension of an arbitrary Kähler Lie algebra (with Kähler form used as cocycle) is turned into a Sasakian algebra [5]. A $3$-dimensional Sasakian Lie algebra is isomorphic to either $\operatorname{\mathfrak{su}}(2)$ or $\operatorname{\mathfrak{sl}}(2,\mathbb{R})$ if simple, or $\mathfrak{aff}(\mathbb{R})\times\mathbb{R}$ or the Heisenberg algebra $\mathfrak{h}_{3}$ if it has centre [25, 16]. The concern then starts with $\dim\mathfrak{g}=5$. In that dimension a centreless Sasakian Lie algebra must be $\operatorname{\mathfrak{su}}(2)\times\mathfrak{aff}(\mathbb{R})$, $\operatorname{\mathfrak{sl}}(2,\mathbb{R})\times\mathfrak{aff}(\mathbb{R})$, or a non-unimodular solvable algebra of the form $\mathbb{R}^{2}\ltimes\mathfrak{h}_{3}$ [5].
Under the assumption that $\mathfrak{g}$ has non-trivial centre, by the way, Cartan’s criterion ensures that solvability is preserved back and forth: $\mathfrak{g}$ solvable $\iff$ $\mathfrak{h}$ solvable.
We recall that if a simply connected Lie group admits lattices then it must be unimodular. Being able to check unimodularity using the trace of the adjoint representation is clearly much better than doing so at the group’s level, if only because the Haar measure might not be known explicitly. Nilpotent, reductive and compact Lie algebras/groups are unimodular, but solvable ones may not ($\mathfrak{aff}(\mathbb{R})$ being the simplest instance). Finding simply connected solvable Lie groups admitting lattices is not an easy task, and few general criteria are known to determine whether a given solvable Lie group admits lattices: one such regards nilpotent Lie groups and is due to Malcev, while Bock [11] addresses interesting non-nilpotent classes. Others, thought less practical ones, can be found in the encyclopedic compendium [27, Ch.2, §3.7]. We just mention that the presence of a lattice in a solvable Lie group forces so-called strong unimodularity [24]. The classification of $4$-dimensional unimodular Lie groups, with and without lattices, can be found for instance in [30]. We also recall that geometric structures on solvable Lie groups are interesting even if they do not give rise to compact examples. For instance, $1$-connected Riemannian homogeneous spaces of non-positive curvature can be directly regarded as solvable groups $G$ with an appropriate left-invariant metric, and much is known in the unimodular case [31, 8, 9]. Specifically for Sasakian Lie groups, the reference articles are [3, 2, 17].
3. Einstein Sasakian algebras with centre
We wish to study the curvature properties of Sasakian algebras $\mathfrak{g}$. Suppose $\xi$ is central, so that the bracket on
$\mathfrak{g}=\mathbb{R}\xi\oplus\mathfrak{h}$
boils down to
(3) $x,y]=-2\omega(x,y)\xi+\theta(x,y)$
on the Kähler reduction $(\mathfrak{h},\theta,J,\langle\cdotp,\cdotp\rangle_{\mathfrak{h}},\omega)$.
In the sequel the sub-/super-script h will denote any object (connection, tensor and the like) calculated on $\mathfrak{h}$. The Riemannian connection and the curvature tensors will be defined in terms of $\langle\cdotp,\cdotp\rangle_{\mathfrak{h}}$.
Lemma 3.1.
Let $\mathfrak{g}$ be Sasakian with centre and $\mathfrak{h}$ its Kähler quotient by $\xi$. The respective curvatures are related by
$R_{x,y}z=R^{\mathfrak{h}}_{x,y}z+\omega(y,z)\varphi(x)-\omega(x,z)\varphi(y)-2\omega(x,y)\varphi(z)$
for all $x,y,z\in\mathfrak{h}$.
Proof.
Take any $x,y,z\in\mathfrak{h}$. Implementing the formulas
(4) $\nabla_{x}y=-\omega(x,y)\xi+\nabla^{\mathfrak{h}}_{x}y,\quad\nabla_{\xi}x=-\varphi(x)$
lifted from [6, Section 2], we obtain
$\nabla_{x}\nabla_{y}z=\nabla_{x}(-\omega(y,z)\xi+\nabla^{\mathfrak{h}}_{y}z)=\omega(y,z)\varphi(x)-\omega(x,\nabla^{\mathfrak{h}}_{y}z)\xi+\nabla^{\mathfrak{h}}_{x}\nabla^{\mathfrak{h}}_{y}z.$
Using (3),
$\nabla_{[x,y]}z=\nabla_{-2\omega(x,y)\xi+\theta(x,y)}z=2\omega(x,y)\varphi(z)-\omega(\theta(x,y),z)\xi+\nabla^{\mathfrak{h}}_{\theta(x,y)}z,$
so eventually:
$\begin{split}R_{x,y}z&=\omega(y,z)\varphi(x)-\omega(x,\nabla^{\mathfrak{h}}_{y}z)\xi+\nabla^{\mathfrak{h}}_{x}\nabla^{\mathfrak{h}}_{y}z-\omega(x,z)\varphi(y)+\omega(y,\nabla^{\mathfrak{h}}_{x}z)\xi\\ &\quad-\nabla^{\mathfrak{h}}_{y}\nabla^{\mathfrak{h}}_{x}z-2\omega(x,y)\varphi(z)+\omega(\theta(x,y),z)\xi-\nabla^{\mathfrak{h}}_{\theta(x,y)}z\\ &=R^{\mathfrak{h}}_{x,y}z+\omega(y,z)\varphi(x)-\omega(x,z)\varphi(y)-2\omega(x,y)\varphi(z)-\omega(x,\nabla^{\mathfrak{h}}_{y}z)\xi\\ &\quad+\omega(y,\nabla^{\mathfrak{h}}_{x}z)\xi+\omega(\theta(x,y),z)\xi.\end{split}$
But $\mathfrak{h}$ is Kähler and $\nabla^{\mathfrak{h}}$ is skew and torsion-free, so:
$\begin{split}-\omega(x,\nabla_{y}^{\mathfrak{h}}z)+\omega(y,\nabla_{x}^{\mathfrak{h}}z)&=-\langle x,\nabla_{y}^{\mathfrak{h}}Jz\rangle+\langle y,\nabla^{\mathfrak{h}}_{x}Jz\rangle=\langle\nabla^{\mathfrak{h}}_{y}x,Jz\rangle-\langle\nabla^{\mathfrak{h}}_{x}y,Jz\rangle\\ &=-\langle\theta(x,y),Jz\rangle=-\omega(\theta(x,y),z).\end{split}$
Substituting this in the previous equation proves the claim. ∎
Lemma 3.2.
The Ricci curvatures of $\mathfrak{g}$ and $\mathfrak{h}$ are related by
$\operatorname{Ric}(x,y)=-2\langle x,y\rangle+\operatorname{Ric}^{\mathfrak{h}}(x,y),\qquad x,y\in\mathfrak{h}.$
Proof.
If we use an orthonormal basis $\{e_{i}\}$ on $\mathfrak{h}$,
$\operatorname{Ric}(x,y)=\langle R_{\xi,x}y,\xi\rangle+{\textstyle\sum}_{i}\langle R_{e_{i},x}y,e_{i}\rangle.$
Invoking [6, Lemma 2.2 (iii)] and computations in [5, Section 2], the first summand equals
$\langle R_{\xi,x}y,\xi\rangle=-\langle y,R_{\xi,x}\xi\rangle=-\langle y,\eta(x)\xi-\eta(\xi)x\rangle=\langle x,y\rangle.$
The other term is found by routine calculations and Lemma 3.1:
$\begin{split}{\textstyle\sum}_{i}\langle R_{e_{i},x}y,e_{i}\rangle&={\textstyle\sum}_{i}\langle R^{\mathfrak{h}}_{e_{i},x}y+\omega(x,y)\varphi(e_{i})-\omega(e_{i},y)\varphi(x)-2\omega(e_{i},x)\varphi(y),e_{i}\rangle\\ &=\operatorname{Ric}^{\mathfrak{h}}(x,y)+{\textstyle\sum}_{i}\langle\omega(x,y)\varphi(e_{i})-\omega(e_{i},y)\varphi(x)-2\omega(e_{i},x)\varphi(y),e_{i}\rangle\\ &=\operatorname{Ric}^{\mathfrak{h}}(x,y)-{\textstyle\sum}_{i}\big(\omega(e_{i},y)\langle Jx,e_{i}\rangle+2\omega(e_{i},x)\langle Jy,e_{i}\rangle\big)\\ &=\operatorname{Ric}^{\mathfrak{h}}(x,y)-\langle Jy,Jx\rangle-2\langle Jx,Jy\rangle\\ &=\operatorname{Ric}^{\mathfrak{h}}(x,y)-3\langle x,y\rangle.\end{split}$
Substituting proves the claim. ∎
The next fact explains the perhaps mysterious constant 2-2 that discriminates the $\eta$-Einstein types of Definition 2.3.
Proposition 3.3.
Let $\mathfrak{g}$ be a Sasakian Lie algebra with centre, and $\mathfrak{h}$ the Kähler quotient. Then $\mathfrak{g}$ is $\eta$-Einstein (with constant $\lambda$) if and only if $\mathfrak{h}$ is Kähler-Einstein (with Einstein constant $\lambda+2$).
Proof.
The $\eta$-Einstein equation reads
$\operatorname{Ric}(x,y)=\lambda\langle x,y\rangle,\qquad x,y\in\mathfrak{h},$
and by Lemma 3.2 this happens precisely when
$\operatorname{Ric}^{\mathfrak{h}}(x,y)=(\lambda+2)\langle x,y\rangle,$
which in turn is saying $\mathfrak{h}$ is Kähler-Einstein. ∎
Corollary 3.4.
A unimodular Sasakian algebra with centre is null $\eta$-Einstein.
Proof.
Of course $\mathfrak{g}$ is unimodular if and only if $\mathfrak{h}$ is. But a unimodular Kähler algebra must be flat [29], whence Ricci-flat a fortiori. Proposition 3.2 forces $\operatorname{Ric}(x,y)=-2\langle x,y\rangle$, making $\mathfrak{g}$ null $\eta$-Einstein. ∎
Example 3.5 (null $\eta$-Einstein solvmanifold).
Let $\mathfrak{h}=\mathbb{R}e_{1}\ltimes_{\operatorname{ad}}\mathbb{R}^{5}$ be the $6$-dimensional, flat Kähler Lie algebra spanned by an orthonormal basis $(e_{1},e_{2})\cup(e_{3},e_{4})\cup(e_{5},e_{6})$ and where the derivation
[…]
acts on $e_{1}^{\perp}=\mathbb{R}^{5}$. Thus, we are prescribing the essential brackets on $\mathfrak{h}$ as
$e_{1},e_{3}]=ae_{4},\qquad[e_{1},e_{4}]=-ae_{3},,\qquad[e_{1},e_{5}]=be_{6},\qquad[e_{1},e_{6}]=-be_{5}.$
Set $Je_{1}=e_{2}$, $Je_{3}=e_{4}$$Je_{5}=e_{6}$, so that the fundamental $2$-form is $\omega=-e^{12}-e^{34}-e^{56}$.
Let us build out of the above a Sasakian algebra $\mathbb{R}\xi\oplus\mathfrak{h}$ of dimension $7$ with non-trivial centre. We extend the bracket on $\mathfrak{h}$ using (3): this adds to the brackets above the further relations
$e_{1},e_{2}]=2\xi,\qquad[e_{3},e_{4}]=2\xi,\qquad[e_{5},e_{6}]=2\xi.$
In this way we may write
$\mathfrak{g}=\mathbb{R}e_{1}\ltimes_{\operatorname{ad}}\underbrace{(\mathfrak{h}_{5}\times\mathbb{R}e_{2})}_{\mathfrak{n}}$
where $\mathfrak{h}_{5}$ stands for the Heisenberg Lie algebra spanned by $(e_{3},e_{4},e_{5},e_{6},\xi)$ (making $\mathfrak{g}$ almost nilpotent). On $\mathfrak{g}$ we now let
[…]
act on $\mathfrak{n}$, and exponentiate it to
$\operatorname{exp}(t\operatorname{ad}_{e_{1}}^{\mathfrak{g}})=\left(\begin{smallmatrix}1&0&0&0&0&0\\ 0&\cos(ta)&-\sin(ta)&0&0&0\\ 0&\sin(ta)&\cos(ta)&0&0&0\\ 0&0&0&\cos(tb)&-\sin(tb)&0\\ 0&0&0&\sin(tb)&\cos(tb)&0\\ 2t&0&0&0&0&1\end{smallmatrix}\right)$
The basis $(e_{2},e_{3},e_{4},e_{5},e_{6},\xi)$ of $\mathfrak{n}$ is rational, so by picking $t\in\mathbb{Z}$ and $a,b\in\frac{\pi}{2}\mathbb{Z}$, the above matrix becomes integer-valued. By [23, Lemma 2.9] we conclude that the 1-connected almost nilpotent Lie group $G$ integrating $\mathfrak{g}$ admits a lattice $\Gamma$, and thus the compact quotient $\G\Gamma\backslash$ G is a Sasakian solvmanifold. The latter is null $\eta$-Einstein as $G$ is unimodular.
The example can be easily generalised to any odd dimension greater than 7 by augmenting (5) using skew-symmetric blocks $\begin{pmatrix}0&-x\\ x&0\end{pmatrix}$, $x>0$, along the diagonal.
4. Centreless Sasakian algebras with $\dim\operatorname{Im}\operatorname{ad}_{\xi}=2$
In this section the working assumption is that $\mathfrak{g}$ is a $(2n+1)$-dimensional Sasakian Lie algebra with no centre. The Reeb vector no longer commutes with everything, and it was shown in [5, Corollary 3.12] that there is an orthogonal splitting
$\mathfrak{g}=\operatorname{Ker}\operatorname{ad}_{\xi}\oplus\operatorname{Im}\operatorname{ad}_{\xi}.$
Let us call these summands
$\mathfrak{s}:=\operatorname{Ker}\operatorname{ad}_{\xi}\quad\text{ and }\quad\mathfrak{q}:=\operatorname{Im}\operatorname{ad}_{\xi}.$
The kernel of $\operatorname{ad}_{\xi}$ is a Sasakian Lie algebra in which $\xi$ is central [5, Proposition 3.13]. Hence we may apply the reduction process of the previous section to obtain
$\mathfrak{s}=\mathbb{R}\xi\oplus\mathfrak{h}_{1},$
where $(\mathfrak{h}_{1},\theta)$ is Kähler and
(6) $x,y]=-2\omega(x,y)\xi+\theta(x,y),\quad[x,\xi]=0,\qquad x,y\in\mathfrak{h}_{1}.$
Therefore $\mathfrak{g}$ decomposes orthogonally as vector space:
(7) $\mathfrak{g}=\mathbb{R}\xi\oplus\underbrace{\mathfrak{h}_{1}\oplus\mathfrak{q}}_{\mathfrak{h}}.$
From now on, whenever $\mathfrak{h}_{1}$ is considered as a Lie algebra we should keep in mind that the Lie bracket on it is $\theta$, as explained earlier.
Let us begin to examine the finer geometric structure on the above splitting. The Jacobi identity for $\xi$, $x\in\mathfrak{h}_{1}$ reads
[…]
We shall concentrate on the case
$\dim\mathfrak{q}=2.$
We saw at the beginning that the mapping $\operatorname{ad}_{\xi}$ is skew-symmetric and commutes with $\varphi$, so we can find an orthonormal basis $(u,v)$ of $\mathfrak{q}$} with $\varphi u=v$ such that
[…]
for some non-zero constant $k$. Then (8) says there exist linear forms $\mu,\gamma\in\mathfrak{h}_{1}^{*}$ such that
[…]
To begin with, we claim that
(11) $u,v]=2\xi+H_{0}\quad\text{ for some }\ H_{0}\in\mathfrak{h}_{1}.$
To see this, the Jacobi identity implies $\xi,[u,v]]=0$, so $u,v]\in\mathfrak{s}$ and $u,v]=t\xi+H_{0}$ for $t\in\mathbb{R}$ and some vector $H_{0}\in\mathfrak{h}_{1}$. The Reeb vector’s factor must equal $t=t\eta(\xi)=\eta(t\xi+H_{0})=-\textup{d}\eta(u,v)=-2\langle u,\varphi(v)\rangle=2$.
Secondly, applying the Jacobi identity with $u,v$ and $x\in\mathfrak{h}_{1}$ we obtain
$\begin{split}0&=\big[x,[u,v]\big]+\big[u,[v,x]\big]+\big[v,[x,u]\big]\\ &=[x,2\xi+H_{0}]+[u,\gamma(x)u-\mu(x)v]+[v,\mu(x)u+\gamma(x)v]\\ &=[x,H_{0}]-2\mu(x)[u,v],\end{split}$
so that $2\mu(x)(2\xi+H_{0})=-2\omega(x,H_{0})\xi+\theta(x,H_{0})$ and finally
$\mu(x)=-\frac{1}{2}\omega(x,H_{0})=-\frac{1}{2}\langle x,JH_{0}\rangle,\quad\text{ and }\quad 2\mu(x)H_{0}=\theta(x,H_{0}).$
Finally, again Jacobi on $x,y\in\mathfrak{h}_{1}$ and $u$ tells
$\begin{split}0&=\underaccent{x,y,u}{\mathfrak{S}}\ \big[x,[y,u]\big]\\ &=[x,\mu(y)u+\gamma(y)v]+[y,-\mu(x)u-\gamma(x)v]+[u,-2\omega(x,y)\xi+\theta(x,y)]\\ &=\bigg(\big(2\omega(x,y)k-\gamma\big(\theta(x,y)\big)\bigg)v-\mu\big(\theta(x,y)\big)u\end{split}$
As $u,v$ were chosen to be independent, either coefficient vanishes. Put more plainly,
$\omega=-\frac{1}{2k}\,\textup{d}^{\mathfrak{h}_{1}}\gamma$
is exact and $\mu$ is closed:
$\textup{d}^{\mathfrak{h}_{1}}\mu=0.$
The normality condition does not provide additional information nor constraints. To sum up, we have proved the following.
Proposition 4.1.
Let $\mathfrak{g}$ be a Sasakian Lie algebra with no centre and $\dim\mathfrak{g}\geqslant 5$, written as (7). If $\dim\operatorname{Im}\operatorname{ad}_{\xi}=2$ then
  1. (1)
    $\mathfrak{h}_{1}$ is a Kähler-exact algebra, meaning $\omega=-\dfrac{1}{2k}\,\textup{d}^{\mathfrak{h}_{1}}\gamma$;
  2. (2)
    $\mu\in\mathfrak{h}_{1}^{*}$ is closed: $\mu(x)=-\dfrac{1}{2}\langle x,JH_{0}\rangle$, and further $2\mu(x)H_{0}=\theta(x,H_{0})$.
From $(2)$ it follows that $0=\mu(\theta(x,y))=-\frac{1}{2}\langle\theta(x,y),JH_{0}\rangle$, so $JH_{0}$ is perpendicular to the derived algebra $\theta(\mathfrak{h}_{1},\mathfrak{h}_{1})$. Taking $x=JH_{0}$ gives $\|H_{0}\|^{2}H_{0}=\theta(JH_{0},H_{0})$, so condition $(1)$ forces
$\|H_{0}\|^{2}\gamma(H_{0})=\gamma\big(\theta(JH_{0},H_{0})\big)=-d\gamma(JH_{0},H_{0})=2k\omega(JH_{0},H_{0})=2k\|H_{0}\|^{2}.$
Therefore, provided $H_{0}\neq 0$, we end up determining the coefficient:
(12) $\gamma(H_{0})=-2k$
The above are only necessary conditions, but they explain that we can reverse-engineer the construction, to show the following.
Proposition 4.2.
Let $(\mathfrak{h}_{1},\theta(\cdot,\cdot),\langle\cdotp,\cdotp\rangle)$ be a metric Lie algebra of dimension $2m$ equipped with a Kähler-exact structure $\omega=-\frac{1}{2k}\textup{d}^{\mathfrak{h}_{1}}\gamma$, for some $k\neq 0$, $\gamma\in\mathfrak{h}_{1}^{*}$. Assume there exist an element $H_{0}\in\theta(\mathfrak{h}_{1},\mathfrak{h}_{1})$ obeying the compatibility conditions:
(13) $JH_{0}\in\theta(\mathfrak{h}_{1},\mathfrak{h}_{1})^{\perp}$
(14) $2\mu(x)H_{0}=\theta(x,H_{0})\quad\text{for every }x\in\mathfrak{h}_{1},$
where $\mu\in\mathfrak{h}_{1}^{*}$ is the closed 1-form given by $\mu(x)=-\frac{1}{2}\langle x,JH_{0}\rangle$, $x\in\mathfrak{h}_{1}$.
Call $\mathfrak{s}=\mathbb{R}\xi\oplus\mathfrak{h}_{1}$ the (Sasakian) central extension of $\mathfrak{h}_{1}$ by the cocycle $\omega$, as described in §3. Choose a $2$-plane $\mathfrak{q}$} with orthonormal basis $(u,v)$ and let
$\mathfrak{g}=\mathfrak{s}\oplus{\mathfrak{q}}$
with bracket prescribed by (6), (9), (10) and (11). This makes $\mathfrak{g}$ a Lie algebra of dimension $2m+3$ with $\mathfrak{s}$ as subalgebra.
Moreover, if we extend the structure $(\varphi,\eta,\xi,\langle\cdotp,\cdotp\rangle)$ on $\mathfrak{s}$ to $\mathfrak{g}$ by
$\varphi(u)=v,\qquad\eta({\mathfrak{q}})=0,$
and the rest in the natural way, $(\mathfrak{g},\varphi,\eta,\xi,\langle\cdotp,\cdotp\rangle)$ becomes a Sasakian Lie algebra with no centre.
As a matter of fact, what we have is a full-on characterisation:
Theorem 4.3.
Any $(2n+1)$-dimensional Sasakian Lie algebra $(\mathfrak{g},\varphi,\eta,\xi,\langle\cdotp,\cdotp\rangle)$ with trivial centre and $\dim\operatorname{Im}\operatorname{ad}_{\xi}=2$ arises by the recipe of the previous proposition.
Examples 4.4.
Take the $4$-dimensional algebra $\mathfrak{h}_{1}=\mathfrak{aff}(\mathbb{R})\times\mathfrak{aff}(\mathbb{R})$. Choose an orthonormal basis $(e_{j})$ with brackets
$\big(0,-e^{12},0,-e^{34}\big)$
Then $\omega=\textup{d}(e^{2}+e^{4}):=-\frac{1}{2k}\textup{d}\gamma$ defines a Kähler-exact structure.
[…]
$\big(0,-e^{12},0,-e^{34},k(2e^{2}+2e^{4}+e^{7})e^{6},k(2e^{2}+2e^{4}-e^{7})e^{5},-2(e^{12}+e^{34}+e^{56})\big),$
is a $7$-dimensional centreless Sasakian algebra. It cannot be solvable because $\mathbb{R}e_{7}\oplus{\mathfrak{q}}$ is simple: $\operatorname{\mathfrak{su}}(2)$ if $k>0$ and $\mathfrak{sl}(2,\mathbb{R})$ for $k<0$.
ii) Another possible choice is $H_{0}=e_{2}$, since $JH_{0}\in\theta(\mathfrak{h}_{1},\mathfrak{h}_{1})^{\perp}=\operatorname{span}\{e_{1},e_{3}\}$. This gives $\mu=\frac{1}{2}e^{1}$, and the end result is now slightly more entangled, algebraically:
$\big(0,-e^{12}-e^{56},0,-e^{34},-k(2e^{2}+2e^{4}-e^{7})e^{6}-\tfrac{1}{2}e^{15},k(2e^{2}+2e^{4}+e^{7})e^{5}-\tfrac{1}{2}e^{16},-2(e^{12}+e^{34}+e^{56})\big),$
In contrast to the instance in i), this one is solvable.
These two examples are indicative of a pattern that we shall address next.
4.1. Issues of solvability
We wish to discuss the solvability of a centreless Sasakian Lie algebra $\mathfrak{g}$ of the type characterised in Theorem 4.3. It turns out that the vector $H_{0}$ appearing in (11) plays an important role in the matter.
Proposition 4.5.
A Sasakian Lie algebra $\mathfrak{g}$ with no centre and $\dim\operatorname{Im}\operatorname{ad}_{\xi}=2$ is solvable, if and only if $H_{0}\neq 0$ and $\mathfrak{h}_{1}$ is solvable.
Proof.
$(\Longrightarrow$) As $\mathfrak{g}$ is solvable, the derived algebra $\mathfrak{g},\mathfrak{g}]$ is nilpotent. We argue by contradiction and suppose $H_{0}=0$. Then equation (11) says $\xi\in[\mathfrak{g},\mathfrak{g}]$ and therefore $\operatorname{ad}_{\xi}$ is nilpotent. It is also skew as the Reeb vector is Killing, so it should vanish altogether, against the assumption. This proves $H_{0}$ cannot vanish. But $\mathfrak{s}$ is a subalgebra in $\mathfrak{g}$ so it is solvable. It also is a central extension of $\mathfrak{h}_{1}\cong{\raisebox{1.99997pt}{$\mathfrak{s}$}\left/\raisebox{-1.99997pt}{$\mathbb{R}\xi$}\right.}$, so $\mathfrak{h}_{1}$ must be solvable too.
[…]
$\begin{split}\mathfrak{g}^{(1)}&=\text{span}\{2\xi+H_{0},u,v\}+\mathfrak{s}^{(1)},\\ \mathfrak{s}^{(1)}&=\{-2\omega(x,y)\xi+\theta(x,y)\mid x,y\in\mathfrak{h}_{1}\}.\end{split}$
We have $2\xi+H_{0},{\mathfrak{q}}]=0$ because
[…]
and similarly $2\xi+H_{0},v]=0$. Secondly, $2\xi+H_{0},\mathfrak{s}^{(1)}]=0$, for
[…]
Finally, ${\mathfrak{q}},\mathfrak{s}^{(1)}]=0$ since
[…]
and $v,\mathfrak{s}^{(1)}]=0$ alike.
From all of that we deduce $\mathfrak{g}^{(2)}=[{\mathfrak{q}},{\mathfrak{q}}]+\mathfrak{s}^{(2)}=\text{span}\{2\xi+H_{0}\}+\mathfrak{s}^{(2)}$ and hence $\mathfrak{g}^{(3)}=\mathfrak{s}^{(3)}$. In other words the derived series of $\mathfrak{g}$ and $\mathfrak{s}$ agree from the third step onwards. But $\mathfrak{s}$ is solvable since $\mathfrak{h}_{1}\cong{\raisebox{1.99997pt}{$\mathfrak{s}$}\left/\raisebox{-1.99997pt}{$\mathbb{R}\xi$}\right.}$ is, by hypothesis. ∎
If we choose $H_{0}=0$ in the construction, $\mathfrak{g}$ will therefore have a semisimple part in its Levi decomposition. Indeed,
Lemma 4.6.
Let $\mathfrak{g}$ be a Sasakian Lie algebra with trivial centre and $\dim\operatorname{Im}\operatorname{ad}_{\xi}=2$. If $H_{0}=0$ then $\operatorname{rad}(\mathfrak{g})=\left\{-\dfrac{\gamma(h)}{k}\xi+h\mid h\in\operatorname{rad}(\mathfrak{h}_{1})\right\}.$
Proof.
Let us recall that
$\operatorname{rad}(\mathfrak{g})=[\mathfrak{g},\mathfrak{g}]^{\perp_{\beta}}=\{z\in\mathfrak{g}\mid\beta(z,y)=0\,\text{ for all}\;y\in[\mathfrak{g},\mathfrak{g}]\},$
where $\beta$ is the Cartan-Killing form of $\mathfrak{g}$. Having $H_{0}=0$ implies $\mu=0$ and $u,v]=2\xi$, so by Proposition 4.2
(15) $\mathfrak{g},\mathfrak{g}]=\mathbb{R}\xi\oplus[\mathfrak{h}_{1},\mathfrak{h}_{1}]\oplus{\mathfrak{q}}.$
Let us, for clarity, write the matrices used to compute $\beta$ in the orthonormal frame $(\xi,e_{1},\ldots,e_{2n},u,v)$ of $\mathfrak{g}$, where $(e_{i})$ is an orthonormal basis of $\mathfrak{h}_{1}$. The matrix of $\operatorname{ad}_{\xi}$ is all zero except for the diagonal $\mathfrak{q}$}-block where (9) holds. For any $h\in\mathfrak{h}_{1}$,
$\operatorname{ad}_{h}=\left(\begin{array}[]{c|ccc|cc}0&\cdots&2\langle Jh,e_{i}\rangle&\cdots&0&0\\ \hline\cr 0&&&&0&0\\ \vdots&&\operatorname{ad}^{\mathfrak{h}_{1}}_{h}&&\vdots&\vdots\\ 0&&&&0&0\\ \hline\cr 0&\cdots&0&\cdots&0&-\gamma(h)\\ 0&\cdots&0&\cdots&\gamma(h)&0\end{array}\right).$
Moreover,
$\operatorname{ad}_{u}=\left(\begin{array}[]{c|ccc|cc}0&\cdots&0&\cdots&0&2\\ \hline\cr 0&&&&0&0\\ \vdots&&0&&\vdots&\vdots\\ 0&&&&0&0\\ \hline\cr 0&\cdots&0&\cdots&0&0\\ -k&\cdots&-\gamma(e_{i})&\cdots&0&0\end{array}\right),\quad\operatorname{ad}_{v}=\left(\begin{array}[]{c|ccc|cc}0&\cdots&0&\cdots&-2&0\\ \hline\cr 0&&&&0&0\\ \vdots&&0&&\vdots&\vdots\\ 0&&&&0&0\\ \hline\cr k&\cdots&\gamma(e_{i})&\cdots&0&0\\ 0&\cdots&0&\cdots&0&0\end{array}\right).$
This gives the values of $\beta$ (for $h,\tilde{h}\in\mathfrak{h}_{1}$):
[…]
Pick an element $z=a\xi+bu+cv+h\in\operatorname{rad}(\mathfrak{g})$, where $a,b,c\in\mathbb{R}$, $h\in\mathfrak{h}_{1}$. Using decomposition (15) we obtain
$\begin{split}0=\beta(z,\xi)&=-2ak^{2}-2k\gamma(h)\Longrightarrow\gamma(h)=-ak\\ 0=\beta(z,u)&=-4bk\Longrightarrow b=0\\ 0=\beta(z,v)&=-4ck\Longrightarrow c=0.\end{split}$
Consequently for $h^{\prime}\in[\mathfrak{h}_{1},\mathfrak{h}_{1}]$
$0=\beta(z,h^{\prime})=-2ak\gamma(h^{\prime})+\beta^{\mathfrak{h}_{1}}(h,h^{\prime})-2\gamma(h)\gamma(h^{\prime})=\beta^{\mathfrak{h}_{1}}(h,h^{\prime}),$
which says $h\in\operatorname{rad}(\mathfrak{h}_{1})$. ∎
Theorem 4.7.
Let $\mathfrak{g}$ be a Sasakian Lie algebra with trivial centre and $\dim\operatorname{Im}\operatorname{ad}_{\xi}=2$ such that $\mathfrak{h}_{1}$ is solvable and $H_{0}=0$ (thus $\mathfrak{g}$ is not solvable). Then the Levi decomposition
$\mathfrak{g}\cong(\mathbb{R}\xi\oplus{\mathfrak{q}})\times\operatorname{rad}(\mathfrak{g})$
is a direct product. Moreover, $\operatorname{rad}(\mathfrak{g})$ is isomorphic to $\mathfrak{h}_{1}$ and the semisimple part is isomorphic to either to $\operatorname{\mathfrak{su}}(2)$ or $\mathfrak{sl}(2,\mathbb{R})$.
Proof.
Since $\mathfrak{h}_{1}$ is solvable it equals its radical $\operatorname{rad}(\mathfrak{h}_{1})$, and it follows from Lemma 4.6 that
$\operatorname{rad}(\mathfrak{g})=\left\{-\frac{\gamma(h)}{k}\xi+h\mid h\in\mathfrak{h}_{1}\right\}.$
Let us denote $\mathfrak{b}:=\text{span}\{\xi,u,v\}=\mathbb{R}\xi\oplus{\mathfrak{q}}$. By virtue of (9), it clearly is a subalgebra isomorphic to $\operatorname{\mathfrak{su}}(2)$ or $\mathfrak{sl}(2,\mathbb{R})$, depending on the sign of $k$.
In order to show that the product is direct rather than semi-direct, we will prove that $\mathfrak{b},\operatorname{rad}(\mathfrak{g})]$ is zero. Consider a radical element $a\xi+h$ against the generic element in $\mathfrak{b}$. First, $\operatorname{ad}_{\xi}(\mathfrak{h}_{1})=0$ implies $\xi,\operatorname{rad}(\mathfrak{g})]=0$. Similarly, ${\mathfrak{q}},\operatorname{rad}(\mathfrak{g})]=0$ by bracketing with $u,v$.
To finish the proof pick a unit vector $e_{1}\in\mathfrak{g}$ with dual $e^{1}=\frac{1}{c}\gamma$ for some real number $c>0$. Extending $e_{1}$ to an orthonormal basis $\{e_{1},e_{2},\ldots,e_{2n}\}$ of $\mathfrak{h}_{1}$ we find
$\operatorname{rad}(\mathfrak{g})=\text{span}\left\{-\frac{c}{k}\xi+e_{1},e_{2},\ldots,e_{2n}\right\},$
which is patently isomorphic to $(\mathfrak{h}_{1},\theta)$ since $\mathfrak{h}_{1}\subseteq\operatorname{Ker}\operatorname{ad}_{\xi}$. ∎
Proposition 4.8.
$\mathfrak{g}$ is unimodular if and only if $\operatorname{tr}\operatorname{ad}^{\mathfrak{h}_{1}}_{x}=-2\mu(x)$, for every $x\in\mathfrak{h}_{1}$.
Proof.
If $x\in\mathfrak{h}_{1}$ it follows that
[…]
and
$\begin{split}\operatorname{tr}\operatorname{ad}^{\mathfrak{s}}_{x}=\langle[x,\xi],\xi\rangle+{\textstyle\sum}_{i}\langle[x,e_{i}],e_{i}\rangle&={\textstyle\sum}_{i}\langle\theta(x,e_{i})-2\omega(x,e_{i})\xi,e_{i}\rangle\\ &={\textstyle\sum}_{i}\langle\theta(x,e_{i}),e_{i}\rangle=\operatorname{tr}\operatorname{ad}^{\mathfrak{h}_{1}}_{x}.\end{split}$
Hence $\operatorname{tr}\operatorname{ad}_{x}=\operatorname{tr}\operatorname{ad}^{\mathfrak{h}_{1}}_{x}+2\mu(x)$ for all $x\in\mathfrak{h}_{1}$, from (9). On the other hand, the full adjoint map of $x\in{\mathfrak{q}}$ only has off-diagonal blocks, whence it is trace-free. Moreover, $\operatorname{ad}_{\xi}$ is traceless because of (9) and $\xi,\mathfrak{s}]=0$. All that reduces the unimodularity to the simpler constraint $\operatorname{tr}\operatorname{ad}^{\mathfrak{h}_{1}}_{x}=-2\mu(x)$, for $x\in\mathfrak{h}_{1}$. ∎
Under the previous assumptions,
Corollary 4.9.
$\mathfrak{g}$ cannot be unimodular in case $H_{0}=0$.
Proof.
Choosing $H_{0}$ to be zero means $\mu(x)=0$, so if $\mathfrak{g}$ were unimodular the previous proposition would make $\operatorname{ad}^{\mathfrak{h}_{1}}$ traceless on $\mathfrak{h}_{1}$. That, though, is not possible since Kähler-exact algebras are not unimodular by a result of [20]. ∎
4.2. Ricci tensors
We begin by computing the Ricci tensor of $\mathfrak{g}$ by brute force. Proposition 2.2 furnishes the following summary:
$\xi$ $y$ $u$ $v$
$\nabla_{\xi}$ 0 Jy-Jy (k1)v(k-1)v (1k)u(1-k)u
$\nabla_{x}$ Jx-Jx $\omega(x,y)\xi+\nabla_{x}^{\mathfrak{h}_{1}}y$ $(\mu(Jx)+\gamma(x))v$ $(\gamma(x)+\mu(Jx))u$
$\nabla_{v}$ $u$ $\mu(Jy)u-\mu(y)v$ $\xi-\tfrac{1}{2}H_{0}$ $\tfrac{1}{2}JH_{0}$
$\nabla_{u}$ v-v $\mu(y)u+\mu(Jy)v$ $\tfrac{1}{2}JH_{0}$ $\xi+\tfrac{1}{2}H_{0}$
with $x,y\in\mathfrak{h}_{1}$. Then
$\operatorname{Ric}(x,y)=\operatorname{Ric}^{\mathfrak{h}_{1}}(x,y)-2\langle x,y\rangle+\langle\nabla_{y}^{\mathfrak{h}_{1}}x,JH_{0}\rangle+\frac{1}{2}\langle H_{0},x\rangle\langle H_{0},y\rangle-\frac{1}{2}\langle JH_{0},x\rangle\langle JH_{0},y\rangle.$
Observe that this is indeed symmetric as $JH_{0}$ is orthogonal to $\mathfrak{h}_{1}^{(1)}$, so $\langle JH_{0},\nabla^{\mathfrak{h}_{1}}_{y}x\rangle=\langle JH_{0},\nabla^{\mathfrak{h}_{1}}_{x}y\rangle$. Moreover, using the Koszul formula for the Levi-Civita connection and Proposition 4.1(2) we may simplify the long formula above to the more significant expression
(16) $\operatorname{Ric}(x,y)=\operatorname{Ric}^{\mathfrak{h}_{1}}(x,y)-2\langle x,y\rangle+\frac{1}{2}\langle\theta(Jx,y),H_{0}\rangle.$
Next, on $\mathfrak{q}$}:
$\operatorname{Ric}(u,v)=\frac{1}{2}{\textstyle\sum}_{i}\big(\langle\nabla^{\mathfrak{h}_{1}}_{e_{i}}H_{0},e_{i}\rangle+\langle H_{0},JH_{0}\rangle\big)=\frac{1}{2}{\textstyle\sum}_{i}\langle\nabla^{\mathfrak{h}_{1}}_{e_{i}}H_{0},e_{i}\rangle,$
and (14) tells
${\textstyle\sum}_{i}\langle\nabla^{\mathfrak{h}_{1}}_{e_{i}}H_{0},e_{i}\rangle={\textstyle\sum}_{i}\langle\theta(e_{i},H_{0}),e_{i}\rangle=-{\textstyle\sum}_{i}\langle e_{i},JH_{0}\rangle\langle H_{0},e_{i}\rangle=-\langle JH_{0},H_{0}\rangle=0.$
Similarly, we have
(17) $\operatorname{Ric}(u,u)=-\|H_{0}\|^{2}+\gamma(H_{0})+2(k-1)-\frac{1}{2}{\textstyle\sum}_{i}\langle\nabla^{\mathfrak{h}_{1}}_{e_{i}}JH_{0},e_{i}\rangle.$
The right-hand side provides an expression for $\operatorname{Ric}(v,v)$, too. The other relevant Ricci terms are easy: $\operatorname{Ric}(x,u)=\operatorname{Ric}(x,v)=0$.
With these preparations in place, we can proceed to refine Proposition 3.3.
Proposition 4.10.
If $H_{0}=0$ then
[…]
Proof.
Let us enforce the above computations under the new assumption, to find
$\begin{array}[]{c}\operatorname{Ric}(x,y)=\operatorname{Ric}^{\mathfrak{h}_{1}}(x,y)-2\langle x,y\rangle,\quad\operatorname{Ric}(u,u)=\operatorname{Ric}(v,v)=2(k-1),\\[5.69054pt] \operatorname{Ric}(u,v)=\operatorname{Ric}(x,u)=\operatorname{Ric}(x,v)=0,\end{array}$
which we write more effectively as matrix on $\mathfrak{h}=\mathfrak{h}_{1}\oplus{\mathfrak{q}}$:
$\operatorname{Ric}=\left(\begin{array}[]{c|c}\operatorname{Ric}^{\mathfrak{h}_{1}}-2\operatorname{\operatorname{Id}}&\\ \hline\cr&2(k-1)\operatorname{\operatorname{Id}}\end{array}\right).$
Immediately, $\mathfrak{g}$ is $\eta$-Einstein precisely when $\operatorname{Ric}^{\mathfrak{h}_{1}}(x,y)=2k\langle x,y\rangle$. ∎
There remains to see to the generic, and more difficult, case in which $H_{0}\neq 0$. Before going head-on into that matter, let us observe the following.
Corollary 4.11.
If $\mathfrak{g}$ is $\eta$-Einstein with $H_{0}\neq 0$, the Einstein constant equals
(18) $\lambda=\frac{1}{2}\operatorname{tr}\operatorname{ad}^{\mathfrak{h}_{1}}_{JH_{0}}-\|H_{0}\|^{2}-2.$
Proof.
We pick up from (17) and push the computation further:
$\begin{split}\langle\nabla^{\mathfrak{h}_{1}}_{e_{i}}JH_{0},e_{i}\rangle&=\langle J\nabla^{\mathfrak{h}_{1}}_{e_{i}}H_{0},e_{i}\rangle=-\langle\nabla^{\mathfrak{h}_{1}}_{e_{i}}H_{0},Je_{i}\rangle\\ &=-\frac{1}{2}\big(\langle\theta(e_{i},H_{0}),Je_{i}\rangle+\theta(Je_{i},H_{0}),e_{i}\rangle+\langle\theta(Je_{i},e_{i}),H\rangle\big)\\ &=-\frac{1}{2}\big(-\langle e_{i},JH_{0}\rangle\langle H_{0},Je_{i}\rangle-\langle Je_{i},JH_{0}\rangle\langle H_{0},e_{i}\rangle+\langle\theta(Je_{i},e_{i}),H_{0}\rangle\big)\\ &=-\frac{1}{2}\big(\langle JH_{0},e_{i}\rangle^{2}-\langle H_{0},e_{i}\rangle^{2}+\langle\theta(Je_{i},e_{i}),H_{0}\rangle\big),\end{split}$
and adding up gives us
$\begin{split}{\textstyle\sum}_{i}\langle\nabla^{\mathfrak{h}_{1}}_{e_{i}}JH_{0},e_{i}\rangle&=-\frac{1}{2}\big(\|JH_{0}\|^{2}-\|H_{0}\|^{2}+{\textstyle\sum}_{i}\langle\theta(Je_{i},e_{i}),H_{0}\rangle\big)\\ &=-\frac{1}{2}{\textstyle\sum}_{i}\langle\theta(Je_{i},e_{i}),H_{0}\rangle=(\textup{d}^{*}\omega)(H_{0})-\operatorname{tr}\operatorname{ad}^{\mathfrak{h}_{1}}_{JH_{0}}\end{split}$
in terms of the codifferential on $\mathfrak{h}_{1}$ (see for instance [7, Section 2]). Of course $\textup{d}^{\ast}\omega=0$ since $\mathfrak{h}_{1}$ is Kähler, and so:
(19) ${\textstyle\sum}_{i}\langle\nabla^{\mathfrak{h}_{1}}_{e_{i}}JH_{0},e_{i}\rangle=-\operatorname{tr}\operatorname{ad}^{\mathfrak{h}_{1}}_{JH_{0}}.$
Substituting what we have just found into (17) and using (12) yields
(20) $\operatorname{Ric}(u,u)=\operatorname{Ric}(v,v)=-\|H_{0}\|^{2}-2+\frac{1}{2}\operatorname{tr}\operatorname{ad}^{\mathfrak{h}_{1}}_{JH_{0}}$
and proves the statement. ∎
The punchline at the end of this section is that in either case ($H_{0}=0$ and $H_{0}\neq 0$) we need to study Kähler-Einstein Lie algebras with exact Kähler form. We shall do just that in the subsequent sections, under the additional simplifying assumption that the Sasakian Lie algebra under consideration (hence the Kähler one as well) is solvable. This leads to studying normal $j$-algebras, which we shall pursue next.
5. Normal j-algebras
Definition 5.1.
Let $\mathfrak{d}$ be a completely solvable real Lie algebra, $J\in\operatorname{\operatorname{End}\,}\mathfrak{d}$ an integrable complex structure and $f\in\mathfrak{d}^{*}$ a $1$-form. The triple $(\mathfrak{d},J,f)$ is called a normal $\bm{j}$-algebra if
$f[Jx,Jy]=f[x,y]\quad\text{ and }\quad f[Jx,x]>0$
for all $x,y\in\mathfrak{d}$.
Such a gadget admits a $J$-invariant, positive-definite inner product
$\langle x,y\rangle:=f[Jx,y].$
The fundamental form associated with the Hermitian pair $\big(J,\langle\cdotp,\cdotp\rangle\big)$ is therefore $\omega=-\textup{d}f$. In other words, a normal $j$-algebra is a completely solvable, Kähler-exact Lie algebra. The following structure theorem is part of Piatetskii-Shapiro’s seminal work.
Theorem 5.2 ([33]).
Let $(\mathfrak{d},J,f)$ be a normal $j$-algebra, and write
$\mathfrak{d}=\mathfrak{n}\oplus\mathfrak{a}$
where $\mathfrak{n}=[\mathfrak{d},\mathfrak{d}]$ and $\mathfrak{a}$ is the orthogonal complement. Then
  • $\mathfrak{a}$ is an Abelian subalgebra and $\mathfrak{n}=\displaystyle\bigoplus_{\alpha\in\Delta}\mathfrak{n}_{\alpha}$ splits into root spaces
    $\mathfrak{n}_{\alpha}=\{x\in\mathfrak{n}\mid[h,x]=\alpha(h)x\text{ for all }h\in\mathfrak{a}\}$
    with $\mathfrak{n}_{\alpha},\mathfrak{n}_{\beta}]\subseteq\mathfrak{n}_{\alpha+\beta}$.
  • Let $\epsilon_{1},\ldots,\epsilon_{r}$ indicate the roots whose root spaces $J$ maps into $\mathfrak{a}$. Then $r=\dim\mathfrak{a}\leqslant\dim\mathfrak{n}$. Up to relabelling, furthermore, all roots are of the form
    $\tfrac{1}{2}\epsilon_{k},\;\epsilon_{k},\text{ for }1\leqslant k\leqslant r,\quad\text{or}\quad\tfrac{1}{2}(\epsilon_{i}\pm\epsilon_{j}),\text{ for }1\leqslant i<j\leqslant r$
    (but some of these combinations may not be roots).
  • If $x\in\mathfrak{n}_{(\epsilon_{i}-\epsilon_{j})/2}$ and $y\in\mathfrak{n}_{\epsilon_{j}/2}$, or $y\in\mathfrak{n}_{\epsilon_{j}}$, are non-zero, then $x,y]\neq 0$,
  • $J\mathfrak{n}_{\epsilon_{k}/2}=\mathfrak{n}_{\epsilon_{k}/2}$,   $J\mathfrak{n}_{(\epsilon_{i}+\epsilon_{j})/2}=\mathfrak{n}_{(\epsilon_{i}-\epsilon_{j})/2}$.
This big theorem should be accompanied by D’Atri’s result that the root spaces are pairwise orthogonal [18]. The roots $\epsilon_{1},\ldots,\epsilon_{r}$ are called distinguished and are linearly independent. Moreover, each $\mathfrak{n}_{\epsilon_{k}}$ has dimension one [18], so one may pick generators $x_{k}\in\mathfrak{n}_{\epsilon_{k}}$ such that $\{Jx_{k}\;;\;1\leqslant k\leqslant r\}$ is a basis of $\mathfrak{a}$ satisfying
(21) $\epsilon_{k}(Jx_{i})=0\ (i\neq k),\quad\epsilon_{k}(Jx_{k})\neq 0.$
Even better, we have $Jx_{k},x_{k}]=\epsilon_{k}(Jx_{k})x_{k}$, so $\omega(Jx_{k},x_{k})=f\big([Jx_{k},x_{k}]\big)=\epsilon_{k}(Jx_{k})f(x_{k})$ and therefore
(22) $0<\|x_{k}\|^{2}=\epsilon_{k}(Jx_{k})f(x_{k}).$
Hence $\epsilon_{k}(Jx_{k})$ and $f(x_{k})$ have the same sign. And finally, by changing the generators’ signs if necessary, we may even assume
(23) $\epsilon_{k}(Jx_{k})>0,\quad f(x_{k})>0.$
This observation will become handy later.
From now on we will use the following notation, for $1\leqslant k\leqslant r$ and $1\leqslant i<j\leqslant r$:
(24) $n_{k}:=\dim\mathfrak{n}_{\epsilon_{k}/2}\in 2\mathbb{Z},\quad\mathfrak{n}_{ij}:=\dim\mathfrak{n}_{(\epsilon_{i}-\epsilon_{j})/2}=\dim\mathfrak{n}_{(\epsilon_{i}+\epsilon_{j})/2}.$
We will also write $\Delta^{\prime}:=\Delta\setminus\{\epsilon_{1},\dots,\epsilon_{r}\}$ for the set of non-distinguished roots.
Example 5.3.
Let $\mathfrak{d}$ be the $6$-dimensional completely solvable Lie algebra
$(0,0,-e^{13}+e^{23},-2e^{14}-2e^{36},-2e^{25},-e^{16}-e^{26}-e^{35})$
with orthonormal basis $(e_{j})$. The complex structure $Je_{1}=e_{4},\,Je_{2}=2e_{5},\,Je_{3}=e_{6}$ and the 1-form $f=-\frac{1}{2}e^{4}-\frac{1}{4}e^{5}$ turn it into a normal $j$-algebra. To find the root spaces note that
$\mathfrak{n}=\text{span}\{e_{3},e_{4},e_{5},e_{6}\}\;\text{ and }\;\mathfrak{a}=\text{span}\{e_{1},e_{2}\}.$
This means there are two distinguished roots $\epsilon_{1},\epsilon_{2}$. Observing that
$e_{1},e_{4}]=\epsilon_{1}(e_{1})e_{4}=2e_{4}\qquad[e_{2},e_{5}]=\epsilon_{2}(e_{2})e_{5}=2e_{5},$
allows to discover that $\epsilon_{1}=2e^{1}$ and $\epsilon_{2}=2e^{2}$.
Regarding non-distinguished roots $\alpha$, let us examine what the possibilities are. If $\alpha=\tfrac{1}{2}\epsilon_{1}$, the root space is even-dimensional for being $J$-invariant, so it must be $\mathfrak{n}_{\alpha}=\text{span}\{e_{3},e_{6}\}$. This would force $e_{2},e_{3}]=0$, a contradiction. A similar reasoning shows $\tfrac{1}{2}\epsilon_{2}\notin\Delta^{\prime}$. Therefore a non-distinguished root might only be one of $\tfrac{1}{2}(\epsilon_{1}\pm\epsilon_{2})$, and in either case Proposition 5.2 gives
$\mathfrak{n}=\mathfrak{n}_{(\epsilon_{1}-\epsilon_{2})/2}\oplus\mathfrak{n}_{(\epsilon_{1}+\epsilon_{2})/2}=\mathbb{R}e_{3}\oplus\mathbb{R}e_{6}.$
5.1. Derivations of normal j-algebras
If $(\mathfrak{d},J,\langle\cdotp,\cdotp\rangle)$ is an arbitrary Hermitian Lie algebra we denote the space of unitary derivations by
$\operatorname{Der}_{u}(\mathfrak{d}):=\{D\in\operatorname{Der}(\mathfrak{d})\mid DJ=JD\text{ and }D\text{ skew-adjoint}\}.$
Lemma 5.4.
Let $\mathfrak{d}=\mathfrak{n}\oplus\mathfrak{a}$ be a normal $j$-algebra. Any skew derivation $D$ kills the Abelian part: $D(\mathfrak{a})=\{0\}$ (and hence $D(\mathfrak{d})\subseteq\mathfrak{n}$).
Proof.
For $h\in\mathfrak{a}$ and $0\neq x\in\mathfrak{n}_{\alpha}$ we have $D[h,x]=[Dh,x]+[h,Dx]$.
The commutator ideal $\mathfrak{n}$ is $D$-invariant, and since $D$ is skew it must preserve the orthogonal complement of $\mathfrak{n}$, namely $\mathfrak{a}$. Therefore $Dh\in\mathfrak{a}$. Now,
$\alpha(h)Dx=\alpha(Dh)x+[h,Dx].$
We may split $Dx=\displaystyle\sum_{\beta\in\Delta}(Dx)_{\beta}$ along $\bigoplus_{\beta\in\Delta}\mathfrak{n}_{\beta}$, so that the above condition refines to
$\alpha(h)\sum_{\beta\in\Delta}(Dx)_{\beta}=\alpha(Dh)x+\sum_{\beta\in\Delta}\beta(h)(Dx)_{\beta}.$
If $\beta=\alpha$ the component in $\mathfrak{n}_{\alpha}$ equals $\alpha(h)(Dx)_{\alpha}=\alpha(Dh)x+\alpha(h)(Dx)_{\alpha}$, which implies $\alpha(Dh)x=0$ and so $\alpha(Dh)=0$. This holds for any $\alpha\in\Delta$, and $\Delta$ contains a basis of $\mathfrak{a}^{*}$, forcing $Dh=0$. ∎
Lemma 5.5.
If $\mathfrak{d}$ is a normal $j$-algebra, any unitary derivation kills each distinguished root space:
$\forall\,D\in\operatorname{Der}_{u}(\mathfrak{d}),\quad D(\mathfrak{n}_{\epsilon_{i}})=\{0\},\quad i=1,\ldots,r.$
Proof.
As $J\mathfrak{n}_{\epsilon_{i}}\subset\mathfrak{a}$, it follows from Lemma 5.4 that $DJ(\mathfrak{n}_{\epsilon_{i}})=\{0\}$. But $D$ and $J$ commute so $JD\mathfrak{n}_{\epsilon_{i}}=\{0\}$, ie $D$ kills $\mathfrak{n}_{\epsilon_{i}}$. ∎
In the light of that, unitary derivations might only salvage non-distinguished root spaces. In the following two technical results we analyse this action in detail.
Lemma 5.6.
If $D\in\operatorname{Der}_{u}(\mathfrak{d})$ then $D(\mathfrak{n}_{\epsilon_{i}/2})\subseteq\mathfrak{n}_{\epsilon_{i}/2}$.
Proof.
Pick $u_{i}\in\mathfrak{n}_{\epsilon_{i}/2}$. As $D$ maps it into $\bigoplus_{\alpha}\mathfrak{n}_{\alpha}$, as before we write $Du_{i}=\sum_{\alpha}(Du_{i})_{\alpha}$. For starters, $\langle Du_{i},\mathfrak{n}_{\epsilon_{j}}\rangle=-\langle u_{i},D\mathfrak{n}_{\epsilon_{j}}\rangle=0$ for any $j$, due to Lemma 5.5. Hence all $(Du_{i})_{\mathfrak{n}_{\epsilon_{j}}}$ vanish, and as a consequence $Du_{i}=\sum_{\alpha\in\Delta^{\prime}}(Du_{i})_{\alpha}$.
We claim that $(Du_{i})_{\alpha}=0$ for all $\alpha\in\Delta^{\prime}$, $\alpha\neq\epsilon_{i}/2$. Pick $h_{j}=Jx_{j}\in\mathfrak{a}$, where $\mathfrak{n}_{\epsilon_{j}}=\mathbb{R}x_{j}$ ($i\neq j)$. Then $D[h_{j},u_{i}]=D\left(\tfrac{1}{2}\epsilon_{i}(h_{j})u_{i}\right)=0$ by (21). On the other hand, Lemma 5.4 tells us $D[h_{j},u_{i}]=[Dh_{j},u_{i}]+[h_{j},Du_{i}]=[h_{j},Du_{i}]$, and
$h_{j},Du_{i}]=\left[[h_{j},\sum_{\alpha\in\Delta^{\prime}}(Du_{i})_{\alpha}\right]=\sum_{\alpha\in\Delta^{\prime}}[h_{j},(Du_{i})_{\alpha}].$
Hence $\sum_{\alpha\in\Delta^{\prime}}\alpha(h_{j})(Du_{i})_{\alpha}=0$, which implies
$\alpha(h_{j})(Du_{i})_{\alpha}=0\quad\text{for all}\quad\alpha\in\Delta^{\prime}.$
If $\alpha$ equals $\tfrac{1}{2}\epsilon_{j}$ or $\pm\tfrac{1}{2}(\epsilon_{j}\pm\epsilon_{k})$ (any sign combination) for $k\neq j$, necessarily $(Du_{i})_{\alpha}=0$. But that holds for any $j\neq i$, so the only chance for $(Du_{i})_{\alpha}$ to be non-zero is that $\alpha=\epsilon_{i}/2$, meaning $Du_{i}\in\mathfrak{n}_{\epsilon_{i}/2}$. ∎
Lemma 5.7.
If $D\in\operatorname{Der}_{u}(\mathfrak{d})$ then $D(\mathfrak{n}_{(\epsilon_{i}\pm\epsilon_{j})/2})\subseteq\mathfrak{n}_{(\epsilon_{i}-\epsilon_{j})/2}\oplus\mathfrak{n}_{(\epsilon_{i}+\epsilon_{j})/2}$ for any $i<j$.
Proof.
We will prove the result only for $\tfrac{1}{2}(\epsilon_{i}+\epsilon_{j})$, as the other case is a straightforward consequence of $[D,J]=0$. Let $u_{ij}\in\mathfrak{n}_{(\epsilon_{i}+\epsilon_{j})/2}$ with $i<j$, so as customary
$Du_{ij}=\sum_{\alpha\in\Delta^{\prime}}(Du_{ij})_{\alpha},\quad\text{with}\quad(Du_{ij})_{\alpha}\in\mathfrak{n}_{\alpha},$
because $\langle Du_{ij},\mathfrak{n}_{\epsilon_{k}}\rangle=-\langle u_{ij},D\mathfrak{n}_{\epsilon_{k}}\rangle=0$ by Lemma 5.5. Pick a generator for $\mathfrak{n}_{\epsilon_{k}}=\mathbb{R}x_{k}$ (with $k\neq i,j$) and set $h_{k}=Jx_{k}\in\mathfrak{a}$, so that
$D[h_{k},u_{ij}]=D\left(\frac{1}{2}(\epsilon_{i}+\epsilon_{j})(h_{k})u_{ij}\right)=0$
by virtue of (21). At the same time, Lemma 5.4 guarantees that
$D[h_{k},u_{ij}]=[h_{k},Du_{ij}]=\left[h_{k},\sum_{\alpha\in\Delta^{\prime}}(Du_{ij})_{\alpha}\right]=\sum_{\alpha\in\Delta^{\prime}}\alpha(h_{k})(Du_{ij})_{\alpha}.$
It follows that $\alpha(h_{k})(Du_{ij})_{\alpha}=0$ for all $\alpha\in\Delta^{\prime}$.
If $\alpha=\tfrac{1}{2}\epsilon_{k}$ or $\pm\tfrac{1}{2}(\epsilon_{k}\pm\epsilon_{l})$ with $l\neq k$, then $(Du_{ij})_{\alpha}=0$. But that holds for all $k\neq i,j$, so we only need to verify that $(Du_{ij})_{\alpha}=0$ when $\alpha=\epsilon_{i}/2$ or $\epsilon_{j}/2$. Now, for any $v_{i}\in\mathfrak{n}_{{\epsilon_{i}/2}}$ Lemma 5.6 says $Dv_{i}\in\mathfrak{n}_{\frac{1}{2}\epsilon_{i}}$ and so
$\langle Du_{ij},v_{i}\rangle=-\langle u_{ij},Dv_{i}\rangle=0.$
Consequently $(Du_{ij})_{\epsilon_{i}/2}=0$. Swapping the indices’ roles produces $(Du_{ij})_{\epsilon_{j}/2}=0$ as well. All in all
$Du_{ij}=(Du_{ij})_{(\epsilon_{i}-\epsilon_{j})/2}+(Du_{ij})_{(\epsilon_{i}+\epsilon_{j})/2},$
completing the proof. ∎
For future reference, we summarise the previous four lemmas as follows: unitary derivations of normal $j$-algebras annihilate the Abelian part $\mathfrak{a}$ and all distinguished root spaces $\mathfrak{n}_{\epsilon_{i}}$, while they preserve every $\mathfrak{n}_{\epsilon_{i}/2}$ and the sums $\mathfrak{n}_{(\epsilon_{i}-\epsilon_{j})/2}\oplus\mathfrak{n}_{(\epsilon_{i}+\epsilon_{j})/2}$ (non-distinguished roots).
As we showed earlier on, normal $j$-algebras are special instances of ‘symplectic-exact’ Lie algebras. Hence they are not unimodular by Hano’s theorem [29] (see [20, Proposition 3.4] as well). Examples of non-unimodular, completely solvable, invariant Kähler structures are provided in [26]. Notwithstanding, we can prove that fact directly from definition 5.1.
Proposition 5.8.
Let $\mathfrak{d}=\mathfrak{a}\oplus\bigoplus_{\alpha}\mathfrak{n}_{\alpha}$ be a normal $j$-algebra, and suppose $\{h_{1},\ldots,h_{r}\}$ a basis of $\mathfrak{a}$ with $h_{i}=Jx_{i}$, $\mathfrak{n}_{\epsilon_{i}}=\mathbb{R}x_{i}$. Then $\operatorname{tr}\operatorname{ad}_{h_{i}}\neq 0$.
Proof.
As $\mathfrak{a}$ is Abelian the trace of $\operatorname{ad}_{h_{i}}$ is calculated just by considering an orthonormal basis $(e_{i})$ of $\mathfrak{n}$, and then $\operatorname{tr}\operatorname{ad}_{h_{i}}={\sum}_{j}\langle\operatorname{ad}_{h_{i}}e_{j},e_{j}\rangle$. Suppose the element $e_{j}$ lives in the root space $\mathfrak{n}_{\alpha}$. Then $\langle[h_{i},e_{j}],e_{j}\rangle=\alpha(h_{i})$ will not be zero if $\alpha$ is $\epsilon_{i}$, $\tfrac{1}{2}\epsilon_{i}$, or $\tfrac{1}{2}(\epsilon_{i}\pm\epsilon_{k})$ with $1\leqslant i<k\leqslant n$. To be more precise, we have
$\alpha$ $\epsilon_{i}$ $\tfrac{1}{2}\epsilon_{i}$ $\tfrac{1}{2}(\epsilon_{i}+\epsilon_{k})$ $\tfrac{1}{2}(\epsilon_{i}-\epsilon_{k})$
$\langle[h_{i},e_{\alpha}],e_{\alpha}\rangle$ $\epsilon_{i}(h_{i})$ $\tfrac{1}{2}\epsilon_{i}(h_{i})$ $\tfrac{1}{2}\epsilon_{i}(h_{i})$ $\tfrac{1}{2}\epsilon_{i}(h_{i})$
Note that the values in the last two columns are the same. Equation (23) says
(25) $\operatorname{tr}\operatorname{ad}_{h_{i}}=\sum_{j=1}^{n}\langle[h_{i},e_{j}],e_{j}\rangle=\epsilon_{i}(h_{i})+\frac{1}{2}\epsilon_{i}(h_{i})n_{i}+2\sum_{i<k}\frac{1}{2}\epsilon_{i}(h_{i})n_{ik}>0$
as dimensions are non-negative. ∎
With all of the above in place we are ready to classify normal $j$-algebras of dimension $4$:
Theorem 5.9.
A four-dimensional normal $j$-algebra is isomorphic (as a Lie algebra) to either $\mathfrak{aff}(\mathbb{R})\times\mathfrak{aff}(\mathbb{R})$ or a semi-direct product $\mathbb{R}\ltimes\mathfrak{h}_{3}$. More precisely, it belongs to one of two families of isomorphic Lie algebras:
  • i)
    $\mathfrak{d}_{a,b}:=\operatorname{span}\{h_{1},x_{1}\}\oplus\operatorname{span}\{h_{2},x_{2}\}$, where $Jx_{1}=h_{1}$, $Jx_{2}=h_{2}$ and
    $h_{1},x_{1}]=ax_{1},\quad[h_{2},x_{2}]=bx_{2},\qquad a,b>0;$
  • ii)
    $\mathfrak{d}_{a}:=\mathbb{R}h_{1}\ltimes\operatorname{span}\{x_{1},x_{2},x_{3}\}$, where $Jx_{1}=h_{1},\;Jx_{3}=x_{2}$ and
    $h_{1},x_{1}]=ax_{1},\quad[h_{1},x_{2}]=\frac{a}{2}x_{2},\quad[h_{1},x_{3}]=\frac{a}{2}x_{3},\quad[x_{2},x_{3}]=ax_{1},\qquad a>0.$
The basis in each case is orthonormal.
Proof.
As $\dim\mathfrak{a}\leqslant\dim[\mathfrak{d},\mathfrak{d}]$, the Abelian part can only be one- or two-dimensional, and we shall examine the occurrences separately.
If $\mathfrak{a}$ is a plane there are $2$ distinguished roots $\epsilon_{1},\epsilon_{2}$, and $\mathfrak{n}=\mathfrak{n}_{\epsilon_{1}}\oplus\mathfrak{n}_{\epsilon_{2}}$ is Abelian, by dimensional reasons. Pick orthonormal generators $x_{i}\in\mathfrak{n}_{\epsilon_{i}}$, and define $h_{i}=Jx_{i}$, an orthonormal basis of $\mathfrak{a}$, so $h_{i},x_{j}]=\epsilon_{j}(h_{i})x_{j}$. It follows from (21) that $\epsilon_{j}(h_{i})=0$ when the indices are different. The numbers $a:=\epsilon_{1}(h_{1})$ and $b:=\epsilon_{2}(h_{2})$ can be supposed positive in view of (23). Rescaling the basis of the two affine factors yields the product $\mathfrak{aff}(\mathbb{R})\times\mathfrak{aff}(\mathbb{R})$.
If $\dim\mathfrak{a}=1$ there exists a unique distinguished root $\epsilon$, and by Theorem 5.2 only one non-distinguished root $\frac{1}{2}\epsilon$. Thus $\mathfrak{n}=\mathfrak{n}_{\epsilon}\oplus\mathfrak{n}_{\epsilon/2}$ with $\dim\mathfrak{n}_{\epsilon/2}=2$. Choose an orthonormal basis for $\mathfrak{n}$ such that $\mathfrak{n}_{\epsilon}=\mathbb{R}x_{1}$ and $\mathfrak{n}_{\epsilon/2}=\text{span}\{x_{2},x_{3}\}$ with $x_{2}=Jx_{3}$ (recall $\mathfrak{n}_{\epsilon/2}$ is $J$-invariant by Theorem 5.2). As $\mathfrak{n}_{\epsilon/2},\mathfrak{n}_{\epsilon/2}]\subseteq\mathfrak{n}_{\epsilon}$, we write $x_{2},x_{3}]=bx_{1}$ for some $b\in\mathbb{R}$. Moreover, $x_{1}$ commutes with $x_{2}$ and $x_{3}$ since $\epsilon+\frac{1}{2}\epsilon$ is not a root. Setting $h_{1}:=Jx_{1}\in\mathfrak{a}$, we have the following brackets:
$h_{1},x_{1}]=ax_{1},\quad[h_{1},x_{2}]=\frac{a}{2}x_{2},\quad[h_{1},x_{3}]=\frac{a}{2}x_{3},\quad[x_{2},x_{3}]=bx_{1}$
where $a=\epsilon_{1}(h_{1})>0$. Applying $f$ to the first and last relationships gives $af(x_{1})=\omega(h_{1},x_{1})=1=\omega(x_{2},x_{3})=bf(x_{1})$, hence $a=b$. We may finally adjust the metric to obtain the standard structure equations of $\mathfrak{h}_{3}$ extended by $h_{1}$. ∎
Remark 5.10.
It is not hard to generalise the family $\mathfrak{d}_{a}$ of Theorem 5.9 to produce normal $j$-algebras with one-dimensional Abelian part: for $n\geqslant 1$ and any real number $a>0$, let $\mathfrak{d}_{a}^{n}$ be the $(2n+2)$-dimensional Lie algebra $\mathfrak{d}_{a}^{n}=\mathfrak{a}\oplus\mathfrak{n}$, where
$\mathfrak{a}=\mathbb{R}h_{1},\quad\mathfrak{n}=\text{span}\{x_{1},u_{1},\ldots,u_{2n}\},$
and the Lie bracket is
$h_{1},x_{1}]=ax_{1},\quad[h_{1},u_{j}]=\frac{1}{2}au_{j},\quad 1\leqslant j\leqslant 2n,\qquad[u_{2i-1},u_{2i}]=ax_{1},\quad 1\leqslant i\leqslant n.$
By declaring the basis $\{h_{1},x_{1},u_{1},\ldots,u_{2n}\}$ orthonormal, and then setting $Jx_{1}=h_{1}$ and $Ju_{2i}=u_{2i-1}$ for $1\leqslant i\leqslant n$, this Lie algebra becomes a normal $j$-algebra. It is almost nilpotent, with nilradical isomorphic to the Heisenberg algebra $\mathfrak{h}_{2n+1}$. For $n=1$ we recover the algebra $\mathfrak{d}_{a}$ of the previous theorem.
Furthermore, it is quite easy to check that any normal $j$-algebra with $1$-dimensional Abelian part is equivalent to one of our $\mathfrak{d}_{a}^{n}$.
Proposition 5.11.
Let $(\mathfrak{d},J,f)$ be a normal $j$-algebra. If $\alpha$ is a non-distinguished root then $f(\mathfrak{n}_{\alpha})=\{0\}$, whereas if $\alpha=\epsilon_{i}$ is distinguished $f(\mathfrak{n}_{\epsilon_{i}})\neq\{0\}$ for all $i$.
Proof.
Let $\epsilon_{1},\ldots,\epsilon_{r}$ denote the distinguished roots. A non-distinguished $\alpha$ must be $\tfrac{1}{2}\epsilon_{i}$ or $\tfrac{1}{2}(\epsilon_{i}\pm\epsilon_{j})$. In the latter case $f(\mathfrak{n}_{\alpha})=0$ [18]. If $\alpha=\tfrac{1}{2}\epsilon_{i}$ let us take an element $y\in\mathfrak{n}_{\epsilon_{i}/2}$. If $\mathfrak{n}_{\epsilon_{i}}=\mathbb{R}x_{i}$ then $Jx_{i},y]=\tfrac{1}{2}\epsilon_{i}(Jx_{i})y$. On the other hand,
$\tfrac{1}{2}\epsilon(Jx_{i})f(y)=f\big([Jx_{i},y]\big)=-df(Jx_{i},y)=\omega(Jx_{i},y)=\langle Jx_{i},Jy\rangle=\langle x_{i},y\rangle=0$
since $x_{i},y$ belong to different root spaces. Hence $f(y)=0$ due to (21), implying $f(\mathfrak{n}_{\epsilon_{i}/2})=0$. Finally, for $\mathfrak{n}_{\epsilon_{i}}=\mathbb{R}x_{i}$ it follows from (23) that $f(x_{i})\neq 0$. ∎
5.2. Einstein normal j-algebras
The Einstein equation for normal $j$-algebras $(\mathfrak{d},J,f)$ is encoded into the condition, proved in [19, Section 1], that there exists a constant $C>0$ such that:
$C=\frac{\epsilon_{k}(Jx_{k})}{f(x_{k})}\left(1+\frac{1}{4}n_{k}+\frac{1}{2}\sum_{j<k}n_{jk}+\frac{1}{2}\sum_{j>k}n_{kj}\right)\qquad\text{ for all }\ 1\leqslant k\leqslant r,$
where $r=\dim\mathfrak{a}$ and the numbers $n_{k},\ n_{ij}$ are as in (24). If we choose the $x_{k}$ to be unit vectors, equation (22) forces $\epsilon_{k}(Jx_{k})=\frac{1}{f(x_{k})}$, and the previous equation becomes the slightly simpler, but much more practical condition
(26) $f(x_{k})^{2}=C^{-1}\left(1+\frac{1}{4}n_{k}+\frac{1}{2}\sum_{j<k}n_{jk}+\frac{1}{2}\sum_{j>k}n_{kj}\right),\qquad k=1,\ldots,r.$
Corollary 5.12.
A normal $j$-algebra $\mathfrak{d}=\mathfrak{a}\oplus\mathfrak{n}$ with $\dim\mathfrak{a}=1$ is Einstein.
Proof.
There is a unique distinguished root $\epsilon$. If $\Delta^{\prime}$ is empty then $\mathfrak{d}=\mathfrak{aff}(\mathbb{R})$, which is Einstein. If there do exist non-distinguished roots, by Theorem 5.2 there is actually only one, namely $\tfrac{1}{2}\epsilon$. In the above formula we then have $r=1$, $n_{jk}=n_{kj}=0$ and $\epsilon_{1}(Jx_{1})>0,\,f(x_{1})>0$ by (23), and the Einstein condition reduces to the quantity $\frac{\epsilon_{1}(Jx_{1})}{f(x_{1})}(1+\frac{1}{4}n_{1})$ being positive. ∎
Remark 5.13.
A consequence of Corollary 5.12 is that the normal $j$-algebras $\mathfrak{d}_{a}^{n}$ of Remark 5.10 are Einstein.
As a further example/application let us establish the following.
Proposition 5.14.
The only Einstein normal $j$-algebra of dimension $2n$ with $\dim\mathfrak{a}=\dim\mathfrak{n}$ is the $n$-fold product $\mathfrak{aff}(\mathbb{R})\times\cdots\times\mathfrak{aff}(\mathbb{R})$.
Proof.
By Theorem 5.2 all roots $\{\epsilon_{1},\ldots,\epsilon_{n}\}$ are distinguished, and if each root space $\mathfrak{n}_{\epsilon_{i}}$ is generated by the unit vector $x_{i}$, then $\{x_{1},\dots,x_{n}\}$ is an orthonormal basis of $\mathfrak{n}$, while $\{h_{1},\ldots,h_{n}\}$ is an orthonormal basis of $\mathfrak{a}$, where $h_{i}=Jx_{i}$. Then the bracket will be
$h_{i},x_{i}]=a_{i}x_{i},\quad 1\leqslant i\leqslant n,$
with all $a_{i}=\epsilon_{i}(h_{i})\neq 0$. At the same time $\textup{d}x^{i}=-a_{i}h^{i}\wedge x^{i}$ says that the Kähler form
$\omega:=h^{1}\wedge x^{1}+\cdots+h^{n}\wedge x^{n}=-\textup{d}\bigg(\sum_{i}\frac{1}{a_{i}}x^{i}\bigg)$
is exact. D’Atri’s theorem ensures there is an Einstein metric precisely if $\frac{a_{k}}{f(x_{k})}=a_{k}^{2}=C>0$ for every $k$, that is, for $a_{1}=\cdots=a_{n}>0$. By rescaling, all these numbers can be taken to be $1$. ∎
The next example provides another family of Einstein normal $j$-algebras.
Example 5.15.
Let $(e_{i})$ be an orthonormal basis of
$\mathfrak{d}=\mathfrak{a}\oplus\mathfrak{n}=\text{span}\{e_{1},e_{2}\}\oplus\text{span}\{e_{3},\ldots,e_{8}\}.$
We have two distinguished roots $\epsilon_{1},\epsilon_{2}$, and let us suppose the root spaces are
$\mathfrak{n}_{\epsilon_{1}}=\mathbb{R}e_{3},\quad\mathfrak{n}_{\epsilon_{2}}=\mathbb{R}e_{4},\quad\mathfrak{n}_{(\epsilon_{1}-\epsilon_{2})/2}=\text{span}\{e_{5},e_{6}\},\quad\mathfrak{n}_{(\epsilon_{1}+\epsilon_{2})/2}=\text{span}\{e_{7},e_{8}\},$
with $Je_{3}=e_{1},Je_{4}=e_{2},Je_{5}=e_{7},Je_{6}=e_{8}$. Equation (21) tells that $\epsilon_{1}=ae^{1},\epsilon_{2}=be^{2}$ with $a,b\neq 0$, and Theorem 5.2 imposes the following:
$\begin{array}[]{c}[e_{1},e_{3}]=ae_{3}\quad[e_{1},e_{5}]=\tfrac{1}{2}ae_{5}\quad[e_{1},e_{6}]=\tfrac{1}{2}ae_{6}\quad[e_{1},e_{7}]=\tfrac{1}{2}ae_{7}\quad[e_{1},e_{8}]=\tfrac{1}{2}ae_{8}\\[5.69054pt] {}[e_{2},e_{4}]=be_{4}\quad[e_{2},e_{5}]=-\tfrac{1}{2}be_{5}\quad[e_{2},e_{6}]=-\tfrac{1}{2}be_{6}\quad[e_{2},e_{7}]=\tfrac{1}{2}be_{7}\quad[e_{2},e_{8}]=\tfrac{1}{2}be_{8}\\[5.69054pt] {}[e_{5},e_{7}]=ce_{3}\quad[e_{5},e_{8}]=ge_{3}\quad[e_{6},e_{7}]=he_{3}\quad[e_{6},e_{8}]=ke_{3}.\end{array}$
for some coefficients to be determined. Applying $f$ to $e_{5},e_{8}]=ge_{3}$ gives $gf(e_{3})=f[e_{5},e_{8}]=\omega(e_{5},e_{8})=\langle e_{5},-e_{6}\rangle=0$, so by (23) we obtain $g=0$. By completely similar computations one finds $h=0$ and c=k=ac=k=-a, so eventually $\mathfrak{d}$ has equations
$\big(0,0,a(-e^{13}+e^{57}+e^{68}),-be^{24},\tfrac{1}{2}(be^{2}-ae^{1})e^{5},\tfrac{1}{2}(be^{2}-ae^{1})e^{6},\tfrac{1}{2}(be^{2}-ae^{1})e^{7},-\tfrac{1}{2}(ae^{1}+be^{2})e^{8}\big).$
The Kähler form $\omega=e^{13}-e^{42}-e^{57}-e^{68}$ can be written as $\omega=-\textup{d}f$ with $f=\frac{1}{a}e^{3}+\frac{1}{b}e^{4}$.
All in all, there is a $2$-parameter family of normal $j$-algebras $\mathfrak{d}$ with $r=2,n_{1}=n_{2}=0,n_{12}=2$. If we take $x_{1}=e_{3},x_{2}=e_{4}$ in equation (26), then $\mathfrak{d}$ is Einstein if and only if $f(e_{3})^{2}=f(e_{4})^{2}$. This happens for $a=\pm b$, so we end up with two curves of Einstein normal $j$-algebras within the big family:
$\begin{array}[]{c}\mathfrak{a}\oplus\big(a(e^{57}+e^{68}-e^{13}),-ae^{24},\tfrac{a}{2}(e^{25}-e^{15}),\tfrac{a}{2}(e^{26}-e^{16}),\tfrac{a}{2}(e^{27}-e^{17}),-\tfrac{a}{2}(e^{18}+e^{28})\big),\\[2.84526pt] \mathfrak{a}\oplus\big(a(e^{57}+e^{68}-e^{13}),ae^{24},-\tfrac{a}{2}(e^{25}+e^{15}),-\tfrac{a}{2}(e^{26}+e^{16}),-\tfrac{a}{2}(e^{27}+e^{17}),-\tfrac{a}{2}(e^{18}-e^{28})\big).\end{array}$
Next we show how to produce concrete examples of $\eta$-Einstein Sasakian Lie algebras beginning with certain normal $j$-algebras. We retain the notation of §4 and Remark 5.10.
Theorem 5.16.
Let $\mathfrak{k}$ be a Kähler-Einstein Lie algebra with exact Kähler form and Einstein constant $c<0$, and $\mathfrak{d}_{a}^{n}$ a $(2n+2)$-dimensional normal $j$-algebra as in Remark 5.10, for some $n\in\mathbb{N}$, $a>0$. Set $\mathfrak{h}_{1}=\mathfrak{d}_{a}^{n}\times\mathfrak{k}$. If $H_{0}=-ax_{1}$ (and $k\neq 0$ is arbitrary), the associated centreless Sasakian Lie algebra $\mathfrak{g}$, constructed using Proposition 4.1, is $\eta$-Einstein if and only if
$c^{2}=\frac{1}{2}a^{2}(n+3).$
This structure is negative $\eta$-Einstein with constant $\lambda=-c^{2}-2$.
Proof.
Corollary 5.12 and Remark 5.13 guarantee that the Kähler metric on $\mathfrak{d}_{a}^{n}$ is Einstein. Moreover, it is straightforward to check that the Ricci curvature equals
(27) $\operatorname{Ric}^{\mathfrak{d}_{a}^{n}}=-a^{2}\left(\frac{n}{2}+1\right)\langle\cdotp,\cdotp\rangle_{\mathfrak{d}_{a}^{n}},$
and so $\operatorname{Ric}^{\mathfrak{h}_{1}}$ has block-diagonal form $\operatorname{Ric}^{\mathfrak{d}_{a}^{n}}\oplus\,c\operatorname{\operatorname{Id}}_{2n}$ in terms of the orthogonal sum $\mathfrak{h}_{1}=\mathfrak{d}_{a}^{n}\oplus\mathfrak{k}$. Therefore, equation (16) gives
$\operatorname{Ric}(h_{1},h_{1})=-a^{2}-\frac{1}{2}a^{2}n-2+\frac{1}{2}\langle\theta(Jh_{1},h_{1}),-ax_{1}\rangle=-\frac{1}{2}a^{2}(n+3)-2$
and $\operatorname{Ric}(x_{1},x_{1})=\operatorname{Ric}(h_{1},h_{1})$, as $\operatorname{Ric}(\varphi x,\varphi y)=\operatorname{Ric}(x,y)$ whenever $x,y\perp\xi$. A direct computation gives the same value for $\operatorname{Ric}(u_{i},u_{i})$, $1\leqslant i\leqslant 2n$, and also for $\operatorname{Ric}(u,u)=\operatorname{Ric}(v,v)=-a^{2}-2+\frac{1}{2}-a\operatorname{tr}\operatorname{ad}_{h_{1}}$ with $u,v\in{\mathfrak{q}}$. Taking a unit vector $x\in\mathfrak{k}$, on the other hand, produces
$\operatorname{Ric}(x,x)=-c^{2}-2+\frac{1}{2}\langle\theta(Jx,x),-ax_{1}\rangle=-c^{2}-2.$
So, for $\mathfrak{g}$ to be $\eta$-Einstein the numbers must match, which happens precisely when $2c^{2}=a^{2}(n+3)$. At last, $c^{2}-2<-2$ says the $\eta$-Einstein structure is negative. ∎
6. Modifications
One of the core lessons we have learnt thus far is that a good chunk of the theory of $\eta$-Einstein Sasakian algebras boils down, so to speak, to the understanding of Kähler-exact solvable algebras. This subject has of course its own interest, irrespective of our Sasakian construction on top of it.
In this part we shall review, and elaborate on, the notion of modified Hermitian algebra as presented in [21, 2]. Modifications, besides being a very flexible and practical algebraic tool, are a core aspect of the Gindikin–Vinberg conjecture on homogeneous Kähler manifolds [22].
Definition 6.1.
Let $\big(\mathfrak{h},[\cdot,\cdot],J,\langle\cdotp,\cdotp\rangle\big)$ be a Hermitian Lie algebra. A linear map $\phi:\mathfrak{h}\to\operatorname{Der}_{u}(\mathfrak{h})$ satisfying
$\phi(x),\phi(y)]=0,\qquad\phi[x,y]=0,\qquad\phi\big(\phi(x)y\big)=0$
for all $x,y\in\mathfrak{h}$ is is called modification map.
Purely for notational purposes we shall ‘curry’ a modification $\phi$ and understand it as a family of unitary derivations $\big\{\phi_{x}\in\operatorname{Der}_{u}(\mathfrak{h})\colon\phi_{x}(y):=\phi(x)y,\ x,y\in\mathfrak{h}\big\}$ parametrised by $\mathfrak{h}$ (more prosaically, we just enforce the exponential law).
Define a modified bracket in the following way:
$x,y]_{\phi}:=[x,y]+\phi_{x}y-\phi_{y}x.$
This is indeed a Lie bracket [17, Proposition 5], and we shall indicate the new structure by $\mathfrak{h}_{\phi}:=(\mathfrak{h},[\cdot,\cdot]_{\phi})$, and write $\textup{d}^{\phi}$ for its Chevalley-Eilenberg differential.
This notion starts to be significant from dimension four upwards, since a modification of a 2-dimensional Lie algebra is trivial. Indeed, for $\mathbb{R}^{2}$, take $e_{1},e_{2}$ orthonormal and let $\phi(v)=\left(\begin{smallmatrix}0&-g(v)\\ g(v)&0\end{smallmatrix}\right)$ be a modification, for some linear map $g$. Then $0=\phi(\phi(e_{1})(e_{1}))=\phi(g(e_{1})e_{2})=g(e_{1})\phi(e_{2})$, and similarly $g(e_{2})=0$. The proof for the non Abelian $2$-dimensional algebra, $\mathfrak{aff}(\mathbb{R})$, follows immediately from the definition of modification map.
Henceforth we shall be concerned with solvable Kähler algebras $\mathfrak{s}$. Dorfmeister showed that any modification $\mathfrak{s}_{\phi}$ stays solvable Kähler. He actually proved the following theorem:
Proposition 6.2 ([21]).
Every solvable Kähler Lie algebra $(\mathfrak{s},J)$ is a modification of a completely solvable semi-direct product of a normal $j$-algebra times an Abelian ideal, both of which are $J$-invariant.
Because of this result, our focus can shift to modified normal $j$-algebras. Specifically in dimension four, for instance, we have:
Proposition 6.3.
A 44-dimensional, modified normal $j$-algebra $\mathfrak{s}_{\phi}$ is either isomorphic to
  • i)
    one of the $\mathfrak{d}_{a,b}$ in Theorem 5.9, or
  • ii)
    a modification of one of the $\mathfrak{d}_{a}$ in Theorem 5.9, with structure equations
    $h_{1},x_{1}]=ax_{1},\quad[h_{1},x_{2}]=\frac{1}{2}ax_{2}+cx_{3},\quad[h_{1},x_{3}]=\frac{1}{2}ax_{3}-cx_{2},\quad[x_{2},x_{3}]=x_{1},$
    where the additional parameter $c\in\mathbb{R}$ accounts for the modification.
Proof.
i) Let $\phi$ be a modification of $\mathfrak{d}_{a,b}$. Lemma 5.4 warrants that it suffices to consider $\mathfrak{a},\mathfrak{n}]_{\phi}$. As $\dim\mathfrak{a}=\dim\mathfrak{n}$, by Theorem 5.2 the only two roots $\epsilon_{1},\epsilon_{2}$ are distinguished. Now, $\phi_{\mathfrak{a}}(\mathfrak{n})=0$ and $\phi_{\mathfrak{n}}(\mathfrak{a})=0$ due to Lemmas 5.45.5, resulting in $\mathfrak{a},\mathfrak{n}]_{\phi}=[\mathfrak{a},\mathfrak{n}]$. Hence any modification acts trivially on $\mathfrak{d}_{a,b}$.
ii) In view of Lemmas 5.45.6, the modification acts on a vector $v\in\mathfrak{d}_{a}$ as
$\phi(v)=\left(\begin{smallmatrix}0&0&0&0\\ 0&0&0&0\\ 0&0&0&-g(v)\\ 0&0&g(v)&0\end{smallmatrix}\right)$
for some linear functional $g$. Letting $c=g(h_{1})$, we recover the structure equations claimed. ∎
Proposition 6.4.
Any modification $\phi$ of a normal $j$-algebra $\mathfrak{d}$ does not alter the derived algebra: $\mathfrak{d},\mathfrak{d}]_{\phi}=[\mathfrak{d},\mathfrak{d}]$.
Proof.
$(\subseteq)$ First of all, $\mathfrak{a},\mathfrak{a}]_{\phi}=[\mathfrak{a},\mathfrak{a}]$ (Lemma 5.4) and clearly $\mathfrak{n},\mathfrak{n}]_{\phi}=[\mathfrak{n},\mathfrak{n}]$. Now take $x\in\mathfrak{a}$, $y=\sum_{\alpha\in\Delta}y_{\alpha}\in\mathfrak{n}=\bigoplus_{\alpha\in\Delta}\mathfrak{n}_{\alpha}$, so
$x,y]_{\phi}-[x,y]=\phi_{x}y=\sum_{\alpha\in\Delta}\phi_{x}y_{\alpha}.$
This combination belongs in $\mathfrak{n}$ by Lemmas 5.65.7, proving the claim.
$(\supseteq)$ Let $\alpha=\epsilon_{i}\neq 0$. As $\mathfrak{a},\mathfrak{n}_{\epsilon_{i}}]_{\phi}=[\mathfrak{a},\mathfrak{n}_{\epsilon_{i}}]+\phi_{\mathfrak{a}}\mathfrak{n}_{\epsilon_{i}}-\phi_{\mathfrak{n}_{\epsilon_{i}}}\mathfrak{a}=[\mathfrak{a},\mathfrak{n}_{\epsilon_{i}}]=\epsilon_{i}(\mathfrak{a})\mathfrak{n}_{\epsilon_{i}}$, we deduce $\mathfrak{n}_{\epsilon_{i}}\subseteq[\mathfrak{d},\mathfrak{d}]_{\phi}$ and therefore $\bigoplus\mathfrak{n}_{\epsilon_{i}}\subseteq[\mathfrak{d},\mathfrak{d}]_{\phi}$.
[…]
$J|_{\mathfrak{n}_{\epsilon_{i}/2}}=\left(\begin{smallmatrix}0&-1&&&\\ 1&0&&&\\ &&\ddots&&\\ &&&0&-1\\ &&&1&0\end{smallmatrix}\right)\qquad\phi_{h}|_{\mathfrak{n}_{\epsilon_{i}/2}}=\left(\begin{smallmatrix}0&-\gamma_{1}(h)&&&\\ \gamma_{1}(h)&0&&&\\ &&\ddots&&\\ &&&0&-\gamma_{d}(h)\\ &&&\gamma_{d}(h)&0\end{smallmatrix}\right)$
can be written in block form. This tells us that
$Je_{k}=f_{k},\quad\phi_{h}e_{k}=\gamma_{k}(h)f_{k},\quad\phi_{h}f_{k}=-\gamma_{k}(h)e_{k},$
so $h,e_{k}]_{\phi}=\tfrac{1}{2}\epsilon_{i}(h)e_{k}+\gamma_{k}(h)f_{k}$, $h,f_{k}]_{\phi}=\tfrac{1}{2}\epsilon_{i}(h)f_{k}-\gamma_{k}(h)e_{k}$, or more compactly
[…]
Note that there exists an element $h\in\mathfrak{a}$ with $\epsilon_{i}(h)\neq 0$, and so the above matrix’ determinant $\frac{1}{4}\epsilon_{i}^{2}(h)+\gamma_{k}^{2}(h)$ is different from zero. Therefore $\operatorname{ad}_{h}^{\phi}$ is an isomorphism on $\text{span}\{e_{k},f_{k}\}$, and thus $\mathfrak{n}_{\epsilon_{i}/2}\subseteq\text{span}\{e_{k},f_{k}\}\subseteq\text{Im}\operatorname{ad}_{h}^{\phi}\subseteq[\mathfrak{d},\mathfrak{d}]_{\phi}$.
For the last type of non-distinguished root, observe that $\mathfrak{n}_{ij}:=\mathfrak{n}_{(\epsilon_{i}-\epsilon_{j})/2}\oplus\mathfrak{n}_{(\epsilon_{i}+\epsilon_{j})/2}$ is both $J$- and $\phi$-invariant (Theorem 5.2 and Lemma 5.7). Choose orthonormal bases $\{u_{1},u_{2},\ldots,u_{r}\}$ and $\{v_{1},v_{2},\ldots,v_{r}\}$ in the two spaces, with $Ju_{i}=v_{i}$. Since $\phi_{h},J]=0$ for any $h\in\mathfrak{a}$, and either is skew, the overall basis $\{u_{i},v_{j}\}$ is orthonormal for $\mathfrak{n}_{ij}$ and yields the normal forms
[…]
Write $\mathfrak{n}_{\epsilon_{i}}=\mathbb{R}x_{i}$, and apply the above to $h_{i}=Jx_{i}\in\mathfrak{a}$ to find
$h_{i},u_{k}]_{\phi}=\frac{1}{2}\epsilon_{i}(h_{i})u_{k}+\phi_{h_{i}}u_{k}\qquad[h_{i},v_{k}]_{\phi}=\frac{1}{2}\epsilon_{i}(h_{i})v_{k}+\phi_{h_{i}}v_{k}$
[…]
In consequence of that, $\mathfrak{n}_{ij}\subseteq\text{Im}\operatorname{ad}^{\phi}_{h_{i}}\subseteq[\mathfrak{d},\mathfrak{d}]_{\phi}$ and $\bigoplus\mathfrak{n}_{ij}\subseteq[\mathfrak{d},\mathfrak{d}]_{\phi}$.
To recap, we have shown that the sums ${\textstyle\bigoplus}\mathfrak{n}_{\epsilon_{i}}$, ${\textstyle\bigoplus}\mathfrak{n}_{\epsilon_{i}/2}$ and ${\textstyle\bigoplus}\mathfrak{n}_{ij}$ all live in $\mathfrak{d},\mathfrak{d}]_{\phi}$, so $\mathfrak{d},\mathfrak{d}]=\bigoplus_{\alpha\in\Delta}\mathfrak{n}_{\alpha}\subseteq[\mathfrak{d},\mathfrak{d}]_{\phi}$. ∎
Proposition 6.5.
For any modification $\phi$ of a normal $j$-algebra, $\operatorname{tr}\operatorname{ad}_{h}=\operatorname{tr}\operatorname{ad}^{\phi}_{h}$ for all $h\in\mathfrak{a}$.
Proof.
Choose an orthonormal basis $(e_{i})$, so $\operatorname{tr}\operatorname{ad}_{h}^{\phi}=\sum_{i}\langle[h,e_{i}]_{\phi},e_{i}\rangle$. Lemma 5.4 ensures that $h,e_{i}]_{\phi}=[h,e_{i}]+\phi_{h}e_{i}-\phi_{e_{i}}h=[h,e_{i}]+\phi_{h}e_{i}$, and the last summand does not really count as $\phi_{h}$ is skew. ∎
6.1. Compact quotients of Sasakian Lie groups
As is already patent in definition 6.1 the complex structure $J$ plays no part in defining the Lie-algebraic structure of the modification. We required $\phi$ to be unitary only for the purposes of adapting the resulting structure to the geometry of normal $j$-algebras, which are the main concern. But the concept makes sense on any metric Lie algebra $\big(\mathfrak{g},\langle\cdotp,\cdotp\rangle\big)$, where a modification is a homomorphism
$\phi\colon\mathfrak{g}\longrightarrow\operatorname{Der}(\mathfrak{g})\cap\operatorname{\mathfrak{so}}(\mathfrak{g})$
such that $\phi([\mathfrak{g},\mathfrak{g}])=0$ and $\phi\big(\operatorname{Im}\phi(\mathfrak{g})\big)=0$. Cortés and Hasegawa [17] had already observed this by showing that a unimodular Sasakian Lie algebra is either $\operatorname{\mathfrak{su}}(2)$, $\operatorname{\mathfrak{sl}}(2,\mathbb{R})$ (simple) or a modification $(\mathfrak{h}_{2n+1})_{\phi}$ of the Heisenberg Lie algebra (solvable). It then follows that a unimodular Sasakian Lie algebra of dimension $\geqslant 5$ is necessarily solvable.
Lemma 6.6.
Let $\mathfrak{h}_{2n+1}$ be the $(2n+1)$-dimensional Heisenberg Lie algebra equipped with an inner product $\langle\cdotp,\cdotp\rangle$, and $\phi$ a modification of it. Then $(\mathfrak{h}_{2n+1})_{\phi}$ has non-trivial center.
Proof.
Consider an orthonormal basis $\{x_{1},\ldots,x_{n},y_{1},\ldots,y_{n},z\}$ of $\mathfrak{h}_{2n+1}$ giving brackets $x_{i},y_{i}]=r_{i}z$ for some $r_{i}>0$, for all $i=1,\ldots,n$. The element $z$ spans the centre. Take $x\in(\mathfrak{h}_{2n+1})_{\phi}$, so that $x,z]_{\phi}=[x,z]+\phi_{x}z-\phi_{z}x=\phi_{x}z$. As derivations preserve the derived algebra, we have $\phi_{x}z=cz$ for some $c\in\mathbb{R}$. This, together with the fact that $\phi_{x}$ is skew, tells that $0=\langle\phi_{x}z,z\rangle=\langle cz,z\rangle=c|z|^{2}$. Eventually, $z$ is a central element in $(\mathfrak{h}_{2n+1})_{\phi}$. ∎
Theorem 6.7.
Let $G$ be a Lie group of dimension $\geqslant 5$ with a left-invariant $\eta$-Einstein Sasakian structure. If it admits lattices, then $G$ is solvable and the structure is null $\eta$-Einstein.
Proof.
The existence of lattices in $G$ is known to guarantee unimodularity (see [34] or [31], for instance). The aforementioned Cortés–Hasegawa theorem says the Lie algebra $\mathfrak{g}$ must be a modification of $\mathfrak{h}_{2n+1}$. The previous lemma warrants non-trivial center, and by Proposition 4.8 the structure has to be null. ∎
Immediately then,
Corollary 6.8.
Every $\eta$-Einstein solvmanifold (of dimension at least $5$) is null.
7. Kähler-exact Lie algebras
We will retain all earlier conventions, but now denote Kähler-exact Lie algebras by $\mathfrak{k}$.
Lemma 7.1.
A Kähler-exact Lie algebra has no (non-trivial) $J$-invariant Abelian subalgebra.
Proof.
Assume the Kähler form is the differential of some functional $f\in\mathfrak{k}^{*}$. If there existed a non-trivial $J$-invariant Abelian subalgebra $\mathfrak{a}$, we could pick an element $0\neq x\in\mathfrak{a}$ and then $0=f[x,Jx]=-\textup{d}f(x,Jx)=-\omega(x,Jx)=\|x\|^{2}$, a contradiction. ∎
Proposition 7.2.
Let $(\mathfrak{d},J,f)$ be a normal $j$-algebra. Any modification $\mathfrak{d}_{\phi}$ is a Kähler-exact Lie algebra with fundamental form $\omega=-\textup{d}^{\phi}f$.
Proof.
We shall examine $\textup{d}^{\phi}f(x,y)=-f\big([x,y]+\phi_{x}y-\phi_{y}x\big)$ for all possible $x,y\in\mathfrak{d}=\mathfrak{a}\oplus\mathfrak{n}$. If $x,y\in\mathfrak{a}$ the bracket does not change (Lemma 5.4) and therefore $\textup{d}^{\phi}f=\textup{d}f$. The same goes for $x,y\in\mathfrak{n}$, due to the definition of modifications. The slightly more involved situation is when $x\in\mathfrak{a}$, $y\in\mathfrak{n}$, in which case we have already seen that
$x,y]_{\phi}=[x,y]+\phi_{x}y.$
Using the root decomposition we might as well assume $y\in\mathfrak{n}_{\alpha}$ for some $\alpha\in\Delta$, which may be distinguished ($\alpha=\epsilon_{i}$ for some $1\leqslant i\leqslant r$) or not ($\alpha\in\Delta^{\prime}$). In the former case Lemma 5.5 implies $\phi_{x}y=0$, so again $\textup{d}^{\phi}f=\textup{d}f$. In the latter, $\phi_{x}y\in\bigoplus_{\beta\in\Delta^{\prime}}\mathfrak{n}_{\beta}$ by Lemmas 5.6, 5.7. According to Proposition 5.11 $f$ vanishes on $\mathfrak{n}_{\beta}$ for any such $\beta$, and therefore $f[x,y]_{\phi}=f[x,y]$. This proves the last case and ends the proof. ∎
The good news is that we can reverse the above result.
Theorem 7.3.
A Kähler-exact solvable Lie algebra is a modification of a normal $j$-algebra.
Proof.
Call $\mathfrak{k}$ the Lie algebra in question. We know from Proposition 6.2 that $\mathfrak{k}$ is a modification $\mathfrak{g}_{\phi}$, where $\mathfrak{g}=\mathfrak{d}\ltimes\mathfrak{b}$ is completely solvable Kähler and obtained as semi-direct product of a normal $j$-algebra $\mathfrak{d}$ and an Abelian ideal $\mathfrak{b}$ (both $J$-invariant).
[…]
Proposition 7.4.
Every $4$-dimensional Kähler-exact Lie algebra is solvable, and hence classified by Proposition 6.3.
Proof.
Straightforward consequence of the previous theorem. ∎
We saw in Proposition 4.2 that the construction of Sasakian algebras from solvable Kähler-exact algebras $\mathfrak{h}_{1}$ depends in a crucial way upon how one picks a special element $H_{0}\neq 0$. Clearly, having too much freedom of choice is undesirable for the purposes of a classification. We show next that this is not the case, for the range of choices turns out to be actually finite.
Proposition 7.5.
Let $(\mathfrak{h}_{1},\theta)$ be a Kähler-exact solvable Lie algebra, and suppose $H_{0}\in\mathfrak{h}_{1}$ satisfies Proposition 4.2. Then either $H_{0}=0$, or $H_{0}$ belongs in a finite subset $F\subset\theta(\mathfrak{h}_{1},\mathfrak{h}_{1})$ of cardinality $|F|=\dim\mathfrak{h}_{1}/\theta(\mathfrak{h}_{1},\mathfrak{h}_{1})$.
Proof.
The algebra $\mathfrak{h}_{1}$ is obtained modifying a normal $j$-algebra $(\mathfrak{d},[\cdot,\cdot])$ via a certain $\phi$, and any non-null $H_{0}$ will live in the derived space $\theta(\mathfrak{h}_{1},\mathfrak{h}_{1})$, hence in the nilradical $\mathfrak{n}=[\mathfrak{d},\mathfrak{d}]$ (Proposition 6.4). As $JH_{0}\in\mathfrak{a}=\mathfrak{n}^{\perp}$, Theorem 5.2 says that $H_{0}\in\bigoplus_{i=1}^{r}\mathfrak{n}_{\epsilon_{i}}$, where $r=\dim\mathfrak{a}$ and $\epsilon_{1},\ldots,\epsilon_{r}$ are the distinguished roots. Since $\theta$ equals the modified bracket $\cdot,\cdot]_{\phi}$, equation (14) reads $2\mu(x)H_{0}=[x,H_{0}]+\phi_{x}H_{0}-\phi_{H_{0}}x$. But the last two summands vanish ($H_{0}\in\mathfrak{n}$ and Lemma 5.5), so we are left with $2\mu(x)H_{0}=[x,H_{0}]$ for all $x\in\mathfrak{h}_{1}$. That means $2\mu$ is a distinguished root $\epsilon_{i}$. Let us then write $H_{0}=tx_{i}$ for a unit generator $x_{i}\in\mathfrak{n}_{\epsilon_{i}}$ and $0\neq t\in\mathbb{R}$. Since $2\mu$ is the metric dual to $JH_{0}$, it follows that $\epsilon_{i}(x)=-t\langle x,h_{i}\rangle$ with $h_{i}=Jx_{i}$. Taking $x=h_{i}$ now produces $t=-\epsilon_{i}(h_{i})$. Altogether $H_{0}$ belongs in the finite set
$\{-\epsilon_{i}(h_{i})x_{i}\mid i=1,\ldots,r\}$
of cardinality $r=\dim\mathfrak{a}=\dim\mathfrak{h}_{1}-\dim\theta(\mathfrak{h}_{1},\mathfrak{h}_{1})$. ∎
7.1. Curvature
It is rather straightforward to check that a general modification $\mathfrak{h}_{\phi}$ has Levi-Civita connection
$\nabla^{\phi}_{x}=\nabla_{x}+\phi_{x}$
(see [17, Proposition 6] as well).
Proposition 7.6.
Modifications preserve the entire Riemann tensor, and hence we have $\operatorname{Ric}^{\phi}=\operatorname{Ric}$ too.
Proof.
Take any $x,z\in\mathfrak{h}$. Then
$R^{\phi}_{z,x}$ $=\nabla^{\phi}_{z}\nabla^{\phi}_{x}-\nabla^{\phi}_{x}\nabla^{\phi}_{z}-\nabla^{\phi}_{[z,x]_{\phi}}$
$=\nabla_{z}\nabla_{x}+\phi_{z}\nabla_{x}+\nabla_{z}\phi_{x}+\phi_{z}\phi_{x}-\nabla_{x}\nabla_{z}-\nabla_{x}\phi_{z}-\phi_{x}\phi_{z}$
$\quad-\nabla_{[z,x]}-\phi_{[z,x]}-\nabla_{\phi_{z}x}-\phi_{\phi_{z}x}+\nabla_{\phi_{x}z}+\phi_{\phi_{x}z}$
$=R_{z,x}+\phi_{z}\nabla_{x}+\nabla_{z}\phi_{x}-\phi_{x}\nabla_{z}-\nabla_{x}\phi_{z}-\nabla_{\phi_{z}x}+\nabla_{\phi_{x}z}$
But any skew derivation $D$, such as our maps $\phi_{x}$, satisfies $\nabla_{x}D+\nabla_{Dx}=D\nabla_{x}$. This means that in the last line above only $R_{z,x}$ survives whilst the other terms cancel each other out. ∎
Let us go back to the setting of §4, where $\mathfrak{g}$ was a centreless Sasakian Lie algebra obtained from a Kähler-exact Lie algebra $(\mathfrak{h}_{1},\theta)$ and some vector $H_{0}\neq 0$. Assuming $\mathfrak{h}_{1}$ is solvable, it follows from Theorem 7.3 that $\mathfrak{h}_{1}$ is a modification $\mathfrak{d}_{\phi}$ of some normal $j$-algebra $\mathfrak{d}$. Using this data, in formula (16) we may replace the term $\operatorname{Ric}^{\mathfrak{h}_{1}}$ by $\operatorname{Ric}^{\mathfrak{d}}$ (Proposition 7.6), and the summand $\langle\theta(Jx,y),H_{0}\rangle$ reads
$\langle\theta(Jx,y),H_{0}\rangle=\langle[Jx,y]+\phi_{Jx}y-\phi_{y}Jx,H_{0}\rangle=\langle[Jx,y],H_{0}\rangle+\langle\phi_{Jx}y,H_{0}\rangle-\langle\phi_{y}Jx,H_{0}\rangle.$
In the right-hand side, the last two terms vanish because, for instance, $\langle\phi_{Jx}y,H_{0}\rangle=-\langle y,\phi_{Jx}H_{0}\rangle=0$ (Lemma 5.5, $H_{0}\in\mathfrak{n}_{\epsilon_{i}}$). Therefore the Ricci tensor reads
$\operatorname{Ric}(x,y)=\operatorname{Ric}^{\mathfrak{d}}(x,y)-2\langle x,y\rangle+\frac{1}{2}\langle[Jx,y],H_{0}\rangle,\qquad x,y\in\mathfrak{h}_{1}.$
The fallout is that when studying Einstein-like conditions, it is enough to consider normal $j$-algebras only.
With that in place we now assume $\mathfrak{h}_{1}$ is a normal $j$-algebra, and therefore all of §5 applies. We have seen that $H_{0}$, if non-zero, is of the form $\epsilon_{i}(h_{i})x_{i}\in\mathfrak{n}_{\epsilon_{i}}$ for some $i=1,\ldots,r$ (Proposition 7.5). The following result exhibits a much needed necessary condition on $\mathfrak{h}_{1}$ that obstructs the construction of $\eta$-Einstein Sasakian structures.
Proposition 7.7.
Consider the centreless Sasakian solvable algebra $\mathfrak{g}$ built from the normal $j$-algebra $\mathfrak{h}_{1}$ and $0\neq H_{0}\in\mathfrak{n}_{\epsilon_{i}}$. If $\mathfrak{g}$ is $\eta$-Einstein then $n_{ji}=0$ for all $j<i$.
Proof.
As usual, let $\mathfrak{a}$ denote the Abelian subalgebra orthogonal to the commutator $\mathfrak{h}_{1},\mathfrak{h}_{1}]$. We borrow a computation from [18, Section 3] to arrive at
(28) $\operatorname{Ric}^{\mathfrak{h}_{1}}(h,h)=-\sum_{\alpha\in\Delta}\alpha(h)^{2}\dim\mathfrak{n}_{\alpha},\qquad\text{ for any }\ h\in\mathfrak{a}.$
Consider the orthonormal basis $\{h_{1},\ldots,h_{r}\}$ of $\mathfrak{a}$, where $h_{i}=Jx_{i}$ and the unit vector $x_{i}$ generates the distinguished root space $\mathfrak{n}_{\epsilon_{i}}$. Then $\alpha(h_{i})\neq 0$ if and only if $\alpha$ is one among $\epsilon_{i}$, $\frac{1}{2}\epsilon_{i}$, $\frac{1}{2}(\epsilon_{i}\pm\epsilon_{j})$ with $i<j$, or $\frac{1}{2}(\epsilon_{j}\pm\epsilon_{i})$ with $j<i$, and
(29) $\operatorname{Ric}^{\mathfrak{h}_{1}}(h_{i},h_{i})=-c^{2}\left(1+\frac{1}{4}n_{i}+\frac{1}{2}\sum_{i<j}n_{ij}+\frac{1}{2}\sum_{j<i}n_{ji}\right)$
where $c:=\epsilon_{i}(h_{i})$. Using (16) with $H_{0}=-cx_{i}$ we obtain
$\begin{split}\operatorname{Ric}(h_{i},h_{i})&=\operatorname{Ric}^{\mathfrak{h}_{1}}(h_{i},h_{i})-2+\frac{1}{2}\langle\theta(Jh_{i},h_{i}),H_{0}\rangle\\ &=\operatorname{Ric}^{\mathfrak{h}_{1}}(h_{i},h_{i})-2-\frac{1}{2}c^{2}=-c^{2}\left(\frac{3}{2}+\frac{1}{4}n_{i}+\frac{1}{2}\sum_{i<j}n_{ij}+\frac{1}{2}\sum_{j<i}n_{ji}\right)-2.\end{split}$
On the other hand, formulas (17), (19) and (25) together force
$\operatorname{Ric}(u,u)=-c^{2}\left(\frac{3}{2}+\frac{1}{4}n_{i}+\frac{1}{2}\sum_{i<j}n_{ij}\right)-2.$
Eventually, if $\mathfrak{g}$ is $\eta$-Einstein necessarily $\sum_{j<i}n_{ji}=0$. But as the $n_{ji}$ are non-negative, they must all vanish. ∎
8. Centreless Sasakian algebras with $\operatorname{Ker}\operatorname{ad}_{\xi}=\mathbb{R}\xi$
To complete the picture we have dealt with throughout the paper, we consider the extreme case, so to speak – that being when $\operatorname{Im}\operatorname{ad}_{\xi}$ is largest possible (codimension one).
Proposition 8.1.
A Sasakian Lie algebra with no centre and $\operatorname{Ker}\operatorname{ad}_{\xi}=\mathbb{R}\xi$ is three-dimensional and simple (that is, $\operatorname{\mathfrak{su}}(2)$ or $\operatorname{\mathfrak{sl}}(2,\mathbb{R})$).
Proof.
[…]
(30) $x,\varphi x]=-2\omega(x,\varphi x)\xi+\theta(x,\varphi x)=2\|x\|^{2}\xi+\theta(x,\varphi x)$
for any $x\in\mathfrak{h}$, where $\theta$ is the usual $\mathfrak{h}$-component of $\cdot$$\cdot$]. This implies $\xi\in[\mathfrak{g},\mathfrak{g}]$ whence
$\mathfrak{g}=[\mathfrak{g},\mathfrak{g}]$
is perfect. It may be semisimple or not. In the former case $\mathfrak{g}$ must be $\operatorname{\mathfrak{su}}(2)$ or $\operatorname{\mathfrak{sl}}(2,\mathbb{R})$, as these are the only semisimple algebras admitting a contact form [12, Theorem 5].
[…]
$\mathfrak{g},\mathfrak{g}]=[\mathfrak{s},\mathfrak{s}]\oplus\big([\mathfrak{s},\mathfrak{r}]+[\mathfrak{r},\mathfrak{r}]\big)\subseteq\mathfrak{s}\oplus\mathfrak{n}.$
The fact that $\mathfrak{g}$ is perfect implies $\mathfrak{n}=\mathfrak{r}$.
[…]
$2\|x\|^{2}\xi=[x,\varphi x]-\theta(x,\varphi x),$
with $\theta(x,\varphi x)\in\mathfrak{h}$. The right-hand side lives in $\mathfrak{h}$ since $\mathfrak{n}$ is an ideal contained in $\mathfrak{h}$. So, if $x\neq 0$ then $\xi\in\mathfrak{h}$, and at the same time $\xi\perp\mathfrak{h}$, which is a contradiction. Therefore $\mathfrak{n}=0$ and $\mathfrak{g}$ is semisimple. ∎
Of course the above proposition can be read off [17] since perfect implies unimodular. Our proof has a flavour of its own and ties in with the earlier discussion.